### v1

# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

Abstract
To optimize large Transformer model training, efficient parallel computing and advanced data management are essential. However, current methods often assume a stable and uniform training workload, neglecting imbalances in data sampling and packing that can impede performance. Specifically, data sampling imbalance arises from uneven sequence length distribution of the training data, while data packing imbalance stems from the discrepancy between the linear memory complexity and quadratic time complexity of the attention mechanism. To address these imbalance issues, we develop Hydraulis, which jointly optimizes the parallel strategies and data assignment. For one thing, we introduce large model training with dynamic heterogeneous parallel strategies in response to the sequence length variations within and across training iterations. For another, we devise a two-stage data assignment approach, which strikes a good balance in terms of the training workloads both within and across model replicas. Empirical results demonstrate that Hydraulis outperforms existing systems by 1.32-2.66$\times$.
1 Introduction
In recent years, large Transformer models [80] have become pivotal in fields such as natural language processing [21, 9, 70], computer vision [10, 23, 56], video processing [7, 2, 26], and speech recognition [22, 4, 30]. Their success relies on pre-training models with vast parameters on extensive datasets [43, 25]. However, this process is both time-consuming and costly, making the acceleration of large model training a pressing concern for both academia and industry.
To move towards Artificial General Intelligence (AGI), input data are condensed into sequences of tokens to achieve a unified representation [60, 23], and models are growing larger in size [43]. Thus, large model training takes as input variable-length sequences and has substantial memory consumption. To address this, approaches from two perspectives have been developed. (a) From a parallel computing perspective, methods like data parallel [52, 87, 88, 5], tensor parallel [77, 62, 46], pipeline parallel [35, 44, 61, 49, 69, 33], sequence and context parallel [36, 54, 51, 8, 29, 27] have been proposed, each with trade-offs in latency and memory usage, and are often combined for better performance. (b) From a data management perspective, data packing techniques are widely used in the training of both language [47, 66, 6, 83, 48] and vision [20, 15] models to minimize redundant computations by combining sequences for simultaneous processing.
Despite these advancements, existing training systems typically leverage these two perspectives in a decoupled manner. (a) In parallel computing, most systems overlook the sequence length variation and adopt a static, homogeneous parallel strategy [77, 72, 91, 38], which deploys all training pipelines with the same parallel scheme throughout the training. [^1] In this work, we use the term “training pipeline” (or “pipeline”) to denote a model replica that can carry out forward and backward propagation individually (i.e., a data parallel group). We use “parallel scheme” to denote the parallel configuration of a pipeline (each pipeline does not necessarily employ any parallel methods, i.e., it can be deployed on only one GPU). Given the available GPUs, a “parallel strategy” describes the collection of parallel schemes of all pipelines deployed. (b) In data management, existing works usually specify a maximum sequence length (a.k.a. context length), pack the original sequences into multiple packed ones, and distribute the packed sequences evenly across machines [66, 48].
In short, current works enforce a static, homogeneous parallel strategy and a fixed-length packing on variable-length data. Nevertheless, as investigated in §3, such a decoupled design fails to take account of two key imbalances.
Firstly, sequences in the real world vary in length by nature, following a long-tail distribution in general [19]. This leads to the data sampling imbalance, manifesting as variable sequence lengths within iterations and inconsistent maximum lengths across iterations. When employing a homogeneous parallel strategy, we must ensure the parallel scheme is memory-saving enough (e.g., with a high tensor parallel degree) to support the maximum sequence length, which compromises the efficiency of short sequences.
The second is the data packing imbalance. Due to the attention mechanism in Transformer, there is a discrepancy between the quadratic time complexity and the linear memory complexity w.r.t. the sequence length [80, 18, 66]. However, existing works adopt fixed-length packing, achieving memory balance but leaving the training workloads imbalanced both within each pipeline and among different pipelines.
Motivated by this, this work manages to address these imbalance issues by jointly optimizing the parallel and packing strategies, putting forward co-designs in both the parallel computing and data management perspectives. To this end, we introduce Hydraulis, an innovative system designed for efficient training of large Transformer models. Our major contributions are outlined as follows.
(I) Training with dynamic heterogeneous parallel strategies. Unlike conventional training that relies on a static, homogeneous parallel strategy, we employ dynamic, heterogeneous strategies to address the data sampling imbalance. Each heterogeneous strategy can deploy pipelines with varying parallel schemes to handle sequences of varying lengths. Divergent strategies can be used for different iterations, allowing each iteration to employ the most suitable strategy for the ad hoc sequence length distribution.
(II) Disaggregating optimization and propagation with subgraphs. We present a novel disaggregation of the optimization and propagation phases in distributed training, which disentangles the divergent model partitioning brought by dynamic heterogeneous strategies. We further introduce the subgraph abstraction to unify the representation of sophisticated communication needs during propagation (for activations) and between optimization and propagation (for parameters and gradients). This allows for flexible combinations of parallel schemes and seamless transitions between parallel strategies.
(III) Two-stage sequence assignment. Given a mini-batch of sequences and an arbitrary strategy, we devise a two-stage management of sequences to eliminate the data packing imbalance. In particular, we formulate an optimization problem that aims to determine how packed sequences can be balanced within each pipeline. Based on this, we also strike a good balance across different pipelines by carefully dispatching the original (unpacked) sequences. Besides, this process also estimates latency for each possible strategy, allowing us to evaluate and select the best one for each iteration.
(IV) Data distribution-aware strategy proposal. Guided by the sequence assignment and for practical considerations, we further propose preparing the candidate strategies that are promising for the given training task in advance, in order to accelerate the strategy enumeration and selection for each iteration. It takes the overall sequence length distribution of the entire dataset into account, and proposes the candidate strategies via a dynamic programming process.
(V) Evaluation. Extensive experiments demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over state-of-the-art training systems, while also offering better scalability.
2 Preliminaries
2.1 Data Sampling and Data Packing
During the training of a Transformer model, each iteration samples a mini-batch of sequences with a pre-defined threshold on the total number of tokens, performs forward and backward propagation on the mini-batch, and updates the model. Besides, the model is typically trained with a pre-defined context length, representing the maximum number of tokens it can process in a single input [68]. Any sequence exceeding this limit is typically truncated to fit within the context length. But regardless of whether the original sequences are truncated or not, the lengths of the sampled sequences will inevitably exhibit variance.
This variability introduces challenges when constructing batches, as sequences need to be aligned to a common length for efficient batch processing. Traditionally, this issue is addressed through data padding, where shorter sequences are padded with special tokens (e.g., [PAD] tokens) to match the length of the longest one [21, 80]. Although padding ensures uniformity in sequence length and is straightforward to implement, it leads to unnecessary computation on the padded tokens, results in wasted computational resources.
Figure 1: Left: Illustration of padding and packing. Right: Comparison of the attention latency during LLaMA2 7B training on Nvidia A800 GPUs with tensor parallel (TP). Standard attention scales quadratically, while varlen achieves linear scaling relative to the sequence length (1K) before packing.
(a) Standard training workflow.
(b) Disaggregate optimization and propagation.
(c) Subgraphs instantiation.
Figure 2: "O, P, G" represents optimizer states, parameters and gradients, respectively. "AG, RS, S, R" represents all-gather, reduce-scatter, send and receive, respectively. "To" represents the data type transfer operator (from 32- to 16-bit floating points), and "LNK" represents the link operator designed by us. (a) The overall workflow of the model states sharding technique combined with mixed-precision training. (b) Our proposed optimization-propagation disaggregation technique, which decouples the distributed strategies for the optimization phase and propagation phase, allowing for arbitrary heterogeneous parallel strategies during propagation. (c) The instantiation of Pull Parameter and Communicate Activation subgraphs during forward propagation. This involves inserting custom operators (e.g., send-receive, all-gather, LNK) to support arbitrary propagation strategies and optimization-propagation interactions.
As shown in Figure 1, an alternative to padding is data packing [47], which seeks to maximize computational efficiency by grouping sequences of varying lengths into a single longer sequence, thereby eliminating the need for padded tokens. To prevent cross-contamination between the packed sequences, positional embeddings and the attention mask must be adjusted to preserve the integrity of the attention mechanism [6, 83, 48]. Fortunately, modern techniques such as FlashAttention v2 have integrated support for block diagonal attention, which natively handles variable-length sequences (i.e., varlen) [18, 17]. Moreover, the varlen interface ensures that the attention complexity scales as $\sum{S_{i}}^{2}$, rather than $(\sum{S_{i}})^{2}$, where $S_{i}$ represents the length of the $i$-th original sequence being packed [66]. This can significantly reduce the computational overhead when using data packing, making it a more practical and efficient approach.
2.2 Parallel Methods, Schemes, and Strategies
As models grow in both parameter size and computational complexity, distributed training across multiple devices becomes essential. Below we introduce popular parallel methods for large Transformer model training.
Data parallel. In data parallel (DP), the model is replicated across multiple devices, with each device processing a different subset of the input data [52, 87, 88, 5]. After computing the gradients locally, devices must synchronize their model updates, typically through a gradient reduction process, to ensure consistency across devices.
Parallel scheme of one pipeline. In this work, we call each data parallel group one training pipeline, and the parallel scheme describes how a pipeline is deployed. We can broadly classify parallel methods for a pipeline into inter- and intra-op parallel, respectively.
The Inter-op parallel is mainly implemented with pipeline parallel (PP). Particularly, PP divides the model into sequential stages, with each stage assigned to a different device. Data from a mini-batch is further split into micro-batches, which flow through the pipeline in a staged manner. (When training Transformer models, each micro-batch accounts for one packed sequence.) Techniques such as micro-batch scheduling and gradient accumulation are often employed, allowing multiple micro-batches to be processed efficiently [35, 44, 61, 49, 69, 33].
Unlike inter-op parallel, there are several intra-op parallel methods for a pipeline. Tensor parallel (TP) partitions each layer across multiple GPUs and exchanges intermediate results (e.g., via all-gather and reduce-scatter operations) to complete the forward and backward passes [77, 62, 46]. Additionally, sequence parallel and context parallel have been explored to handle extremely long sequences, where a single sequence is partitioned across multiple devices to avoid out-of-memory (OOM) errors [36, 54, 51, 8, 29, 27].
Parallel strategy. Building on the parallel scheme of a single pipeline, the parallel strategy refers to the collection of parallel schemes of all pipelines. However, existing works mainly leverage the same parallel schemes for all pipelines, resulting in a homogeneous parallel strategy. Such a homogeneous design fails to take account of the imbalances caused by data sampling and data packing, as discussed in §3.
2.3 Model States Sharding
To reduce memory consumption during training, model states sharding techniques (a.k.a. ZeRO [71] and FSDP [89, 88]) are widely used. These methods extend standard data parallel (DP) by partitioning model states (i.e., parameters, gradients, and optimizer states) across multiple devices, rather than replicating them across all data parallel groups (i.e., pipelines).
Figure 3: The dot plot on the left shows the fluctuation of maximum sequence length, emphasizing inter-iteration imbalance, while the bar chart on the right presents the distribution of sequence lengths within an iteration (20th iteration), highlighting intra-iteration imbalance.
The sharding technique is combined with mixed-precision training [59, 65], which uses both 16-bit floating point (FP16 or BF16) and 32-bit floating point (FP32) precisions to balance efficiency and numerical stability. Figure 2(a) illustrates an example with each pipeline deployed with a tensor parallel degree of 2. The 32-bit optimizer states and parameters (which consume more memory but are critical for ensuring convergence) are sharded across GPUs, while the 16-bit parameters and gradients (which are more memory-efficient and computationally faster) are maintained during propagation.
As aforementioned, current works mainly adopt a homogeneous parallel strategy, so they usually align the sharding with the strategy, allowing distributed state transformations to be easily achieved via all-gather and reduce-scatter operations. However, since such transformations can overlap with computation, it is also feasible to adopt more nuanced, non-aligned parameter pulling and gradient pushing using customized communication strategies, which will be elaborated in §5.2.
3 Imbalances Analysis
In this section, we analyze the imbalance issues encountered during data sampling and packing in training on real-world datasets, and discuss potential optimization opportunities. To ease the analysis, we exemplify the imbalance issues by training the LLaMA 7B model with the CommonCrawl [78] and GitHub [16] datasets over 8 GPUs. The context length is 32K and the batch size is set as 100K for each iteration.
3.1 Data Sampling Imbalance
We first concentrate on the imbalance in terms of data sampling. As shown in Figure 3, we observe two key issues: (a) within iterations, sequence lengths are unevenly distributed, with more short sequences and fewer long ones; (b) across iterations, the sequence distribution varies significantly, and the maximum sequence length fluctuates considerably, often falling short of the context length. In the following, we analyze the two issues, respectively.
Intra-iteration imbalance. The uneven sequence length distribution within an iteration can be attributed to the inherent property of data. In particular, sequences in the real world generally follow a long-tail distribution, where short sequences dominate and long sequences are rare. Existing studies have also observed such a distribution in many popular training datasets [19]. This imbalance poses both challenges and opportunities when it comes to optimizing overall performance.
The challenges stem from the trade-offs between memory usage and throughput offered by different numbers of GPUs and parallel schemes. As shown in Figure 4, increasing the number of GPUs generally leads to a lower throughput per GPU but allows for a longer maximum supportable sequence length. Meanwhile, under the same number of GPUs, different parallel schemes (e.g., PP, TP and their combinations) also give divergent trade-offs. As aforementioned in §2.2, existing works utilize all GPUs in the cluster uniformly with a homogeneous parallel strategy tailored for the longest context length, compromising the training efficiency for memory.
Nevertheless, opportunities arise when adapting the trade-offs to sequences of varying lengths: short sequences, which are more frequent, can be processed using faster but more memory-intensive parallel schemes, while the fewer long sequences can be assigned to slower yet more memory-efficient parallel schemes. Therefore, adopting heterogeneous parallel strategies— where non-unique parallel schemes are applied on different proportions of GPUs based on different sequence lengths— enables better performance (§5).
Figure 4: Memory and throughput trade-offs. "OOM" indicates out-of-memory, and "N/A" indicates not available. The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs. Throughput is measured in tokens per second.
Inter-iteration imbalance. While intra-iteration imbalance arises from the distribution of sequence lengths within a single mini-batch, another challenge comes from the distribution variations across mini-batches. Particularly, due to the long-tail distribution, rare long sequences may not appear in every iteration (even with a large batch size, the maximum sequence length still varies, as detailed in Appendix A.1). This leads to fluctuations in data distribution and maximum sequence lengths across iterations. However, current works overlook this issue, using a static parallel strategy for all iterations, which is unsuitable. For example, when an iteration consists mainly of short sequences, parallel schemes designed for long sequences become overkill. To address this, we propose to dynamically adjust the parallel strategy based on the characteristics of each mini-batch to better align the workload, thereby enhancing overall performance across iterations (§5).
Figure 5: The numbers of original sequences (classified by their lengths) in different micro-batches after packing (each micro-batch accounts for one packed sequence).
3.2 Data Packing Imbalance
We then focus on the imbalance issues of data packing. As shown in Figure 5, we observe that due to varying sequence lengths, there are significant differences in the distributions of original sequences within the packed sequences, both among different micro-batches within the same pipeline and across different pipelines. In the following, we will analyze the impact of each aspect individually.
Intra-pipeline imbalance. Since the memory consumption of each micro-batch is linear to the total length after packing, existing works commonly adopt fixed-length packing methods, like bin packing [86, 66], to balance memory usage. However, due to the quadratic time complexity of attention [80, 18, 66], the computation latency can hardly be balanced with fixed-length packing— for example with Figure 1 again, attention latency can be nearly 20×20$\times$20 × smaller when packing 32 sequences of 1K compared to a single 32K sequence.
In pipeline parallel (PP), this leads to uneven execution times across micro-batches, yet common pipeline scheduling, such as 1F1B [33] or zero-bubble [69], requires uniform execution time of micro-batches to minimize pipeline bubbles. This highlights the importance of taking account of both the linear memory complexity and quadratic time complexity, rather than considering solely the memory as in fixed-length packing. To address this, we devise a nuanced packing strategy that considers the variation in original sequence lengths, striking a balance in terms of the execution time across different micro-batches within the memory constraint (§6.1).
Inter-pipeline imbalance. Although different pipelines execute concurrently, the training latency is dictated by the slowest pipeline due to the need for gradients synchronization. Simply packing sequences to a fixed length has a similar issue of unbalanced computation load. In some cases, if the sequence number after packing cannot be evenly distributed across pipelines, a more severe imbalance will arise. In contrast to current systems that first pack sequences and then attempt to distribute them evenly, we suggest reversing this process. Specifically, our work first dispatches sequences of varying lengths to different pipelines, and then focuses on packing them within each pipeline. This innovative dispatching paradigm ensures a more balanced load across pipelines, before organizing sequences into micro-batches (§6.2).
4 Overview
[…]
We develop Hydraulis to address the training imbalances described above. Figure 6 illustrates the overview.
To tackle the data sampling imbalance, we develop a brand new system that supports training with dynamic heterogeneous parallel strategies (§5). In each heterogeneous parallel strategy, the training pipelines can be associated with non-unique parallel schemes to differentiate between long and short sequences for each iteration, addressing the intra-iteration imbalance. And we further support adjusting the strategies dynamically across different iterations to tackle the inter-iteration imbalance.
To tackle the data packing imbalance, given a mini-batch of sequences and a strategy, we devise a two-stage assignment (§6). The first stage determines how to dispatch these sequences of varying lengths to different pipelines to solve the inter-pipeline imbalance. Then, the second stage meticulously packs the sequences within each pipeline to form different micro-batches, addressing the intra-pipeline imbalance. This two-stage assignment process also estimates the latency of the given strategy to train the given mini-batch.
[…]
5 System Designs
In this section, we start by mathematically defining our heterogeneous parallel strategies, then introduce the system techniques proposed to support these strategies and seamless transitions between them.
5.1 Dynamic Heterogeneous Strategies
[…]
Next, we define our heterogeneous strategy space as the linear span of the parallel scheme space: $\mathbb{S}=\text{span}(\mathbb{P})$, meaning that each heterogeneous strategy can be represented as a linear combination of parallel schemes: $S=\sum_{i=1}^{K}d_{i}\times P_{i}$, which indicates that $S$ consists of $\sum_{i=1}^{K}d_{i}$ pipelines of $K$ different parallel schemes, with $d_{i}$ pipelines deployed with $P_{i}$.
Strategy execution and transition. During each training iteration, given a mini-batch of sequences, we select the most suitable $S$ to execute. [^2] The selection in each iteration is based on the two-stage sequence assignment (§6), and we propose the candidate strategies by analyzing the sequence length distribution of the entire dataset (§7). The pipelines will be assigned with non-overlapping sequences and execute the forward and backward propagation concurrently, which is termed the propagation phase. Then, akin to traditional data parallel (DP), all pipelines synchronize the model gradients and perform updates, which is termed the optimization phase. Since different iterations inevitably select non-unique strategies, strategy transitions between iterations are necessary.
System challenges. However, enabling the system to support various pipelines, each configured with a different $P$, is complex due to the need to manage diverse communication patterns across devices. Furthermore, dynamically transitioning between strategies at the iteration level presents additional difficulties, as the model states mappings on devices can vary significantly across different $S$. To address these issues, we develop an optimization-propagation disaggregation approach (§5.2) and a subgraph-based abstraction (§5.3). These system designs successfully meet our requirements of training with dynamic, heterogeneous strategies.
5.2 Optimization-propagation Disaggregation
As illustrated in Figure 2(b), unlike traditional methods that employ a uniform distributed training strategy, we separate the optimization phase from the propagation phase, decoupling the GPU’s role as a distributed storage from its computational role during propagation. Specifically, this approach maintains a constant partitioning for FP32 parameters and optimizer states, while allowing any parallel strategy to be used during propagation, enabling the system to support dynamic heterogeneous strategies and seamless transitions between them. [^3] We use a fully sharded strategy in optimization phase, yet any other zero-redundancy sharding strategies are also feasible. When propagation requires access to FP32 parameters and optimizer states, we perform pull and push interactions, which correspond to the retrieval of parameters and the synchronization and scattering of gradients, respectively. These interactions occur only at the beginning and end of a training iteration, referred to the steps 1-2 and steps 4-5 of Figure 2(b).
To effectively manage pull and push interactions, we develop a specialized LNK (link) operator to handle complex interaction patterns. As depicted on the right side of Figure 7, this operator utilizes three core primitives: send, receive, and local movement, centered around a fundamental unit called the mutual slice, which represents the common partitioning granularity when model shardings interact between the propagation and optimization phases. Given that the optimization states are partitioned on each GPU without redundancy, the pull and push schemes can be uniquely determined using these three primitives on the non-duplicate mutual slices.
In particular, assume there are $N$ GPUs, the optimization phase employs a fully sharded strategy, and the propagation phase uses a strategy $S=\sum_{i=1}^{K}d_{i}\times P_{i}$ with $D=\sum_{i=1}^{K}d_{i}$ pipelines. Let $\overline{TP}_{\text{max}}=\max_{i\in[1,K]}\texttt{TP}(P_{i})$. For the pull interaction, the mutual slices are created by evenly partitioning the global parameter into $\mathtt{LCM}\left(N,\overline{TP}_{\text{max}}\right)$ segments (where $\mathtt{LCM}$ denotes the least common multiple). Each GPU identifies its necessary mutual slices. If a slice is local, it uses the local movement operation; otherwise, it employs send and receive to exchange slices. Similarly, for the push interaction, as shown on the left side of Figure 7, a reduce-scatter is first performed at the corresponding location to synchronize the gradient. Then the partitioning granularity of the mutual slice is $\mathtt{LCM}\left(N,D\times\overline{TP}_{\text{max}}\right)$, with the additional $D$ accounting for the gradient’s reduce-scatter. The LNK can be established still using the send, receive and local movement primitives.
Figure 7: Diagram of RS (reduce-scatter) and LNK (link) operations during push. Pink indicates the gradient, blue the optimizer states, and green the mutual slice. The optimizer states are partitioned with a granularity of 4, and the gradient after reduce-scatter has a granularity of 6. Thus, the mutual slice has a granularity of 12 (i.e., the least common multiple of 4 and 6).
Compared to training with homogeneous strategies, introducing the LNK operator incurs additional overhead to resolve the discrepancy between strategies. However, the send and receive communication can be overlapped with computation during propagation, and the local movement incurs only minimal latency, so the overhead is small (as evaluated in §8.4). Last but not least, our design offers great generalization capabilities— when the optimization and propagation use the same strategy, the LNK operator in the pull will be translated into an all-gather operation, and the LNK operator in the push will be pruned (leaving a reduce-scatter operation).
5.3 Subgraphs
To tackle the challenge of representing diverse heterogeneous strategies and their complex optimization-propagation interactions, we introduce subgraphs— subsets of the complete computation graph. The key observation is that different strategies only have unique patterns in some specific operators, while most user-defined computational operators remain unchanged. Therefore, subgraphs are developed to encapsulate these unique operators from different strategies for efficient representation, management and scheduling.
Figure 8: Scheduling subgraphs with pipeline parallel (PP). In the first PP stage, Pull Parameter subgraphs are scheduled immediately after the last operators requiring the parameter, and Push Gradient subgraphs follow right after the operators that produce or accumulate the gradient. In subsequent PP stages, Pull Parameter and Push Gradient are grouped at the beginning and end, respectively, to utilize inherent PP bubbles. This scheduling scheme effectively hides communication latency within the Pull Parameter and Push Gradient subgraphs.
In particular, we define three subgraph types: Pull Parameter, Push Gradient, and Communicate Activation. Operators for interactions between optimization and propagation phases, like LNK, are placed in the first two, while activation-related communication operators (e.g., all-gather and reduce-scatter for TP, send-receive for PP) are included in the third.
Subgraphs instantiation. As illustrated at the bottom of Figure 2(c), users start by defining various computational operators to describe the training model. We then identify all potential points in the computational graph where different heterogeneous strategies might intervene and preserve several empty subgraphs (e.g., between transformer modules for potential pipeline parallel, before and after matrix multiplication for potential tensor parallel). Given a strategy $S$, the user-defined operators are directly integrated into a new, strategy-specific graph. We then determine how to instantiate all predefined subgraphs with appropriate operators. For example, as depicted at the top of Figure 2(c), a data transfer operator and a strategy-specific LNK operator are added to instantiate a Pull Parameter subgraph. This instantiation process ultimately forms a complete graph for a specific $S$, enabling different GPUs to extract and execute their respective operators.
Subgraphs scheduling. During execution, we optimize overlap by controlling the execution order of subgraphs and computational operators. Each subgraph can be treated as a cohesive unit and integrated into the scheduler as a fused operator. As illustrated in Figure 8, by using a typical cudaStream to execute the subgraphs and interleaving them with computational operators, we enable communication within subgraphs to overlap with computation, thereby reducing latency.
6 Two-stage Sequence Assignment
In this section, we present our two-stage sequence assignment approach. Specifically, given a mini-batch of sequences and an arbitrary strategy, we establish an optimization problem to pack sequences precisely, and ensure balanced sequence dispatching based on load estimates of each pipeline, addressing both intra- and inter-pipeline imbalances. In each iteration, we apply the assignment to all candidate strategies, estimate the propagation latency, and select the optimal one.
6.1 Sequence Packing within Pipeline
We first introduce the packing strategy within each pipeline. Assume there are $U$ sequences with lengths $\{l_{i}\}_{i=1}^{U}$ assigned to a pipeline configured with parallel scheme $P$. To pack these sequences into multiple micro-batches, there are two essential terms to determine: the number of micro-batches to craft (denoted as $V$) and how each sequence should be assigned.
Cost Model. For a given $P$, there is a maximum sequence length it can support, denoted as $\texttt{MaxLen}(P)$ (i.e., sequence exceeding this length will cause OOM on $P$), which can be determined in advance using our memory model (detailed in Appendix B.1). Additionally, we introduce a latency model, $\texttt{T}(l,P)$, which estimates the time taken to run a sequence of length $l$ on $P$ for a single forward and backward propagation of a single pipeline stage (detailed in Appendix B.2).
Solving for one $V$. The problem is easy to solve if the number of micro-batches $V$ is given. To be specific, as mentioned in §2.1, by employing the varlen functionality of FlashAttention v2 [18, 17], the latency of a packed sequence can be approximated as the sum of the latency of the individual original sequences. Meanwhile, to address the intra-pipeline imbalance, our goal is to minimize the duration of the longest-running micro-batch. Thus, we can formulate the following problem on top of the minimum sum partition (MSP) problem:
$\operatorname*{arg\,min}_{o_{i,j}}$ $\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o_{i,j}\times\texttt{T}\left(l_{i},P\right)}\right)\times(\texttt{PP}(P)-1+V)}$ (1)
s.t. $\sum\limits_{1\leq j\leq V}{o_{i,j}}=1,o_{i,j}\in\left\{0,1\right\},\forall i\in\left[1,U\right]$
$\sum\limits_{1\leq i\leq U}{o_{i,j}\times l_{i}}\leq\texttt{MaxLen}\left(P\right),\forall j\in\left[1,V\right]$
where $o_{i,j}\in\{0,1\}$ indicates whether the $i$-th sequence is packed into the $j$-th micro-batch. Compared to the MSP problem, the problem defined in Eq. (1) additionally ensures that the length of each packed sequence must not exceed $\texttt{MaxLen}(P)$, and we multiply the term $\texttt{PP}(P)-1+V$ to represent the propagation latency of the entire pipeline.
Enumeration of $V$. To locate the optimal number of micro-batches $V^{*}$, we can simply enumerate the value for $V\in[1,U]$, solve Eq. (1), and choose the best one. Note that Eq. (1) is an Integer Linear Programming (ILP) problem, but its complexity is only related to the number of sequences assigned to $P$ (i.e., $U$) and the number of micro-batches (i.e., $V$). These do not increase with cluster size, and the enumeration can be done in parallel, ensuring fast problem-solving even on large-scale clusters (detailed in Appendix E). Additionally, we can prune unnecessary enumerations, where a too-small $V$ results in a high bubble rate and a too-large $V$ leads to excessive kernel launch overheads. Detailed pruning methods are provided in Appendix C.
6.2 Sequence Dispatching across Pipelines
Next, we consider how the sequences in a mini-batch should be dispatched across the pipelines to address the inter-pipeline imbalance. Consider a specific strategy $S=P_{1}+\ldots+P_{D}$ satisfying $\texttt{MaxLen}\left(P_{1}\right)\geq\ldots\geq\texttt{MaxLen}\left(P_{D}\right)$. Assume a mini-batch contains $B$ sequences with lengths $\{l_{i}\}_{i=1}^{B}$. We denote $J_{i}$ as the number of pipelines that the $i$-th sequence can run on, satisfying $\texttt{MaxLen}\left(P_{J_{i}}\right)\geq l_{i}>\texttt{MaxLen}\left(P_{J_{i}+1}\right)$. Thus, $l_{i}$ can be dispatched to any of $P_{1},\ldots,P_{J_{i}}$. Since the pipelines execute concurrently, our goal is to dispatch the $B$ sequences across the $D$ pipelines while minimizing the propagation latency of the longest-running pipeline.
However, the propagation latency of each pipeline cannot be directly computed without solving the packing within it (i.e., §6.1). Fortunately, it is feasible to estimate the lower bound of latency. In particular, let $m_{i,j}\in\{0,1\}$ indicate whether the $i$-th sequence is assigned to $P_{j}$, then the lower bound of running time of the $j$-th pipeline can be expressed as (detailed derivation in Appendix D.1):
$\small\sum_{i=1}^{B}m_{i,j}\times\texttt{T}(l_{i},P_{j})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$ (2)
Denote Eq. (2) as $\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$, then the dispatching can be formulated as a (0,1)-ILP problem:
$\operatorname*{arg\,min}_{m_{i,j}}$ $\max\limits_{1\leq j\leq D}\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$ (3)
s.t. $\sum\limits_{1\leq j\leq J_{i}}{m_{i,j}}=1,\forall i\in\left[1,B\right]$
$m_{i,j}\in\left\{0,1\right\},\forall i\in\left[1,B\right],\forall j\in\left[1,D\right]$
The number of decision variables of Eq. (3) is linear w.r.t. number of sequences in a mini-batch (i.e., $B$) and the number of pipelines in the strategy (i.e., $D$). In practice, as the cluster scale expands, both will increase proportionally, making the assignment problem-solving cost unacceptable in large-scale scenarios. Fortunately, there are extensive studies about how to approximately solve assignment problems in polynomial time with a guaranteed error bound [24, 76, 12]. Motivated as such, we implement an approximate solver to Eq. (3), which utilizes multiple random trials to greedily dispatch sequences to pipelines and ultimately select the best trial. Empirical results demonstrate that the approximate solver achieves nearly the same objectives as directly solving Eq. (3), but with greater speed. More details about our approximate solver can be found in Appendix D and E.
7 Data Distribution-aware Strategy Proposal
Ideally, for each iteration, we could explore all possible strategies, estimate their propagation latency via the two-stage sequence assignment, and find the optimal one. However, given the vast size of the strategy space $\mathbb{S}$, this is impractical. Thus, we analyze the dataset in advance and propose a refined subset $\hat{\mathbb{S}}\subset\mathbb{S}$, which only contains the most promising candidates.
Figure 9: Dynamic programming with non-integer values. Different colors represent various $P$. The segments indicate sequences within each dataset interval, with their lengths corresponding to latency. Rounding the number of $P$ requires scaling the resulting latency.
Problem definition. To address both intra- and inter-iteration imbalances of data sampling, the proposal is based on two considerations. Firstly, since intra-iteration imbalance typically reflects the expected distribution of the dataset, we aim to preserve strategies that demonstrate an overall advantage over the entire dataset. Secondly, due to inter-iteration imbalance, characterized by variations in maximum sequence lengths, we aim to identify non-unique strategies across different maximum lengths. Then, we have the following problem.
Problem 1. Given the number of GPUs $N$, the dataset $\mathbb{D}$, and an arbitrary sequence length $L$. With sequence lengths randomly sampled within $[0,L]$ over $\mathbb{D}$, we seek the optimal heterogeneous strategy $S^{*}_{L|\mathbb{D}}$ from $\mathbb{S}$ that minimizes the expected cost. Then the proposed subset is defined as the union: $\hat{\mathbb{S}}=\bigcup_{L\leq L_{\text{max}}}\{S^{*}_{L|\mathbb{D}}\}$, where $L_{\text{max}}$ is the maximum sequence length to support (i.e., context length).
We first simplify the problem by forcing each pipeline to process only the sequences within a specific length interval, rather than any arbitrary lengths. This allows the problem to be solved via dynamic programming. Later we lift the restriction by continuous relaxation, enhancing the solution.
Dynamic programming. We define $t[n][l]$ as the minimum achievable propagation latency using up to $n$ GPUs to process all sequences in the dataset with lengths not exceeding $l$. Our goal is to determine the value $t[N][L]$.
The initial states are obvious that $t[n][0]=0,\forall n>0$ and $t[0][l]=\infty,\forall l>0$. Next, we focus on the state transition for dynamic programming. Let $\mathbb{D}_{(l-l^{\prime},l]}$ denote the sub-dataset containing sequences within the interval $(l-l^{\prime},l]$, and assume the parallel scheme space comprises $K$ unique schemes $\{P_{k}\}_{k=1}^{K}$. Based on our simplification, we only need to consider assigning this sub-dataset to one kind of scheme. Suppose it is assigned to $d$ pipelines of $P_{k}$ and $P_{k}$ requires $\texttt{N}\left(P_{k}\right)$ GPUs, then this assignment consumes $d\times\texttt{N}\left(P_{k}\right)$ GPUs in total. Besides, we must ensure $\texttt{MaxLen}\left(P_{k}\right)\geq l$. Then the propagation latency of this sub-dataset can be approximated as $\frac{1}{d}\sum_{x\in\mathbb{D}_{(l-l^{\prime},l]}}\texttt{T}(x,P_{k})$. Meanwhile, the minimum propagation latency for the remaining sequences is $t\left[n-d\times\texttt{N}\left(P_{k}\right)\right]\left[l-l^{\prime}\right]$. Therefore, we have the following state transition expression:
$t[n][l]=\min\left(t[n-1][l],\min_{(k,d)\in{Cond}_{n,l},l^{\prime}<l}\left\{{T}_{k,d,l^{\prime}}\right\}\right),\ \textup{where}$ (4)
${T}_{k,d,l^{\prime}}=\max\left(t[n-d\times\texttt{N}\left(P_{k}\right)][l-l^{\prime}],\frac{1}{d}\sum_{x\in\mathbb{D}_{(l-l^{\prime},l]}}\texttt{T}(x,P_{k})\right)$
${Cond}_{n,l}=\left\{(k,d)\mid\texttt{MaxLen}\left(P_{k}\right)\geq l,\,d\times\texttt{N}\left(P_{k}\right)\leq n\right\}$
The use of $\max(\cdot,\cdot)$ in $T_{k,d,l^{\prime}}$ is because the slowest pipeline determines the overall latency. Based on the dynamic programming, we also obtain the corresponding optimal strategy $S[n][l]$ upon the calculation of $t[n][l]$, and $S[N][L]$ will be viewed as the solution $S^{*}_{L|\mathbb{D}}$ under $L$ in Problem 1.
By leveraging the optimal sub-structure property of dynamic programming, once we determine $S[N][L_{\text{max}}]$, we obtain $S[N][L]$ for all $L\leq L_{\text{max}}$. We can then select all unique $S[N][L]$ from $L=1$ to $L=L_{max}$ to form $\hat{\mathbb{S}}$.
Continuous relaxation. Based on the dynamic programming, we lift the restriction that each pipeline can only handle a sequence over a single length interval. Specifically, we execute the dynamic programming with non-integer values for $n$ and $d$. This means sequences in a length interval can use “a portion” of a pipeline, enabling a complete pipeline to support multiple intervals. Consequently, the final solution $S[N][L]$ might involve having non-integer numbers, as exemplified in Figure 9. In such cases, we round and adjust to include all nearby integer solutions in $\hat{\mathbb{S}}$, while ensuring that the total number of GPUs does not exceed $N$ [^4] The whole algorithm’s worst-case time complexity is $O(KN^{2}L_{\text{max}}^{2})$. In our experiments, we set the step size of $n$ and $d$ to 0.1 and $l$ to 128, with the actual time cost detailed in Appendix E.
Undoubtedly, even though the original non-integer solution achieves optimal results, rounding compromises the optimality. Additionally, the data distribution in different iterations does not perfectly align with that of the entire dataset. Due to these limitations, we can only suggest potential strategy candidates, whereas the two-stage sequence assignment is still necessary for selecting optimal strategies.
8 Evaluation
Figure 10: End-to-end evaluation. We test the workloads of three configurations, F1-F3, in each scenario. "Max Seq" represents the number of tokens in the longest sequence, while "Iter" denotes the total number of tokens sampled in a single iteration. Error bars indicate the standard deviation across iterations.
8.1 Experimental Setup
Environment. Our experiments are conducted with 8 GPU nodes, each containing 8 Nvidia A800 GPUs (64 GPUs in total). GPUs within a node are connected via NVLink with a bandwidth of 400 GB/s, and nodes are interconnected using Infini-Band with a bandwidth of 200 GB/s.
Baselines. We compare Hydraulis with Megatron-LM [77, 62, 46], DeepSpeed [72, 36], and HotSPa [28]. Both Megatron-LM and DeepSpeed employ a single, static homogeneous strategy throughout the training process. HotSPa uses heuristics to configure multiple homogeneous strategies and enables switching between them within every iteration. However, due to the overhead of switching, HotSPa is limited to setting 3-4 strategies. We tune the strategies for all competitors to achieve the best performance.
Models and datasets. Most existing large models are based on the decoder-only Transformer architecture. We select the widely used LLaMA2 series models and conduct experiments with 7B, 13B and 32B model sizes [79]. For datasets, we use CommonCrawl [78] and GitHub [16], which are widely used open-source datasets and have served as the major training corpora for many prestigious Transformer models.
Worloads and metrics. By default, we sample 100K tokens per iteration [^5] We consider 100K tokens per iteration for our 64-GPU cluster, scaled proportionally from the 16M used in LLaMA with an 8192-GPU cluster [25]. with a context length of 32K. And the Adam optimizer [45] is used. We conduct 20 warm-up iterations and then summarize the throughput (tokens per second) or per-iteration latency over 100 iterations.
Figure 11: Comparison of scalability across various factors, using the fastest one as the reference for relative changes. Shaded areas indicate the standard deviation across iterations.
(a) Case study w.r.t parallel strategies.
(b) Case study w.r.t sequence assignment.
Figure 12: (a) Latency over 100 training iterations, with colors indicating the strategy selected for each iteration. (b) Comparison of load balancing between Hydraulis and Megatron-LM. We present the maximum and minimum propagation latency among all ranks for 5 iterations.
8.2 End-to-end Experiments
We first evaluate Hydraulis against three baselines across various configurations (baselines’ strategies are detailed in Appendix F). The results in Figure 10 demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over the best baselines.
As DeepSpeed and Megatron-LM are limited to a single fixed strategy, they must adopt less efficient strategies to accommodate the memory consumption of long sequences, degrading the overall efficiency given most sequences are relatively short. Although HotSPa is capable of using multiple homogeneous strategies within an iteration to train sequences of varying lengths, it requires multiple times of model hot switching. This incurs significant overheads, especially when per-iteration time is short (e.g., in the “GitHub 7B 8GPUs F1” scenario, the average iteration time is merely 6.7s, while nearly 2s is spent on hot switching). On the contrary, Hydraulis handles variable-length sequences using dynamic heterogeneous strategies, while eliminating the overhead of strategy transition via disaggregating optimization and propagation.
[…]
8.3 Scalability
We then assess the scalability of different systems by examining three factors: (a) number of GPUs, (b) maximum sequence length, and (c) tokens per iteration. Using the LLaMA 7B model and the CommonCrawl dataset, we start with a default setup of 16 GPUs, a maximum sequence length of 32K, and 200K tokens per iteration. By systematically altering one factor while keeping the others constant, we gain insights into each factor’s impact. The results are presented in Figure 11.
Number of GPUs. As the number of GPUs increases, all systems experience a sub-linear reduction in latency due to the communication overhead introduced by parallelism. Specifically, for HotSPa, the communication overhead from hot switching does not decrease with the number of GPUs, causing its overall performance to be surpassed by Megatron-LM when using 64 GPUs. In contrast, Hydraulis introduces minimal additional communication overhead (only the LNK operators), which can overlap with computation. Consequently, its latency consistently outperforms other baselines.
Maximum sequence length. As the maximum sequence length grows, DeepSpeed and Megatron-LM resort to less efficient strategies, resulting in a significant increase in latency. Notably, when Megatron-LM scales from 64K to 128K, it must enable cross-node TP, which greatly increases latency. In contrast, by customizing parallel strategies for sequences of different lengths, HotSPa and Hydraulis are less sensitive to maximum sequence length changes, with Hydraulis demonstrating superior performance over HotSPa.
Tokens per iteration. As the number of tokens per iteration increases, all systems experience an almost linear increase in latency. However, HotSPa and Hydraulis exhibit a slightly lower growth rate, as they take account of the long-tail distribution of sequence lengths within iterations.
8.4 Performance Interpretation
Figure 13: Ablation study. Error bars indicate the standard deviation across iterations. Total throughput changes as different components of Hydraulis are incrementally integrated.
To further evaluate the impact of the proposed techniques, we conduct experiments by training the 32B model with the CommonCrawl dataset over 64 GPUs.
Case study. Figure 12(a) shows the optimal strategies selected by Hydraulis, along with the corresponding latencies, and summarizes the frequency of different strategies over 100 iterations. It can be seen that: (a) the actual propagation latency closely matches our estimation in most cases, and (b) variations across iterations lead to different strategy choices.
[…]
Additionally, we examine the load balancing. Figure 12(b) shows that Megatron-LM, with fixed-length packing and even dispatching of packed sequences, results in a latency difference of 0.36-0.65$\times$ between the slowest and fastest ranks. In contrast, Hydraulis reduces this gap to just 0.02-0.10$\times$, verifying the significance of the two-stage sequence assignment.
Ablation study. Figure 13 presents an ablation study by progressively enabling different techniques in Hydraulis to evaluate their contributions: (a) Starting with a static, homogeneous parallel strategy, we dispatch sequences of varying lengths according to Eq. (3) and pack them to the maximum length. (b) Instead of packing sequences to the maximum length, we apply packing based on the results of Eq. (1). (c) We replace the original homogeneous parallel strategy with a heterogeneous strategy that supports the maximum sequence length. (d) Finally, we enable dynamic heterogeneous strategies, utilizing all Hydraulis designs.
The results show consistent improvements across (a)-(d), demonstrating that each component of the system plays a critical role in enhancing overall performance.
Latency breakdown. We further investigate the effectiveness of optimization-propagation disaggregation compared to conventional model states sharding, as well as the benefits brought by subgraphs scheduling. As shown in Table 1, we first adopt a fixed homogeneous parallel strategy (with the two-stage sequence assignment enabled), which utilizes the conventional all-gather and reduce-scatter interaction patterns (Figure 2(a)). Then, we employ dynamic heterogeneous strategies, introducing additional LNK operators (Figure 2(b)). Besides, we experiment both with and without subgraphs scheduling to measure the latency improvements from overlapping pull and push interactions with computation.
Homo Hetero w/o Overlap Hetero w/ Overlap
P 6.78s 4.29s (1.58×1.58$\times$1.58 ×) 4.29s (1.58×1.58$\times$1.58 ×)
O 0.02s 0.02s (1.00×1.$0$) 0.02s (1.00×1.$0$)
O-P 1.11s 1.71s (0.69×0.69$\times$0.69 ×) 1.18s (0.94×0.94$\times$0.94 ×)
E2E 7.91s 6.02s (1.31×1.31$\times$1.31 ×) 5.49s (1.44×1.$4$)
Table 1: Breakdown of latency per iteration for different training strategies. "P" denotes propagation (Figure 2(b) step 3), "O" refers to optimization (Figure 2(b) step 6), and "O-P" represents the remaining overhead from the interaction between the two phases (Figure 2(b) steps 1-2 and 4-5). "E2E" stands for end-to-end training latency.
It can be observed that the use of dynamic heterogeneous strategies improves the performance of propagation. However, the interaction between optimization and propagation becomes slower due to the additional communication overhead introduced by LNK operators. Despite this, the overhead remains relatively small and can be efficiently reduced by overlapping with computation, leading to only a minimal increase in latency. Moreover, this additional latency does not scale with the number of tokens per iteration, demonstrating the viability of disaggregating optimization and propagation.
9 Related Work
Optimizations of distributed training. Numerous studies aim to enhance the efficiency of distributed training. In the realm of parallel methods, advancements in pipeline parallel focus on minimizing bubble time through enhanced scheduling [69, 49, 39, 85], while tensor parallel has progressed by overlapping communication with computation [82, 37, 11, 40]. In the context of model states sharding, efforts have been made to balance memory usage and latency through finer-grained sharding [13], efficient communication [81, 57], and multi-level storage solutions [73, 90]. Notably, these approaches are compatible with ours. There are also efforts focusing on automatically searching for the optimal parallel strategies [91, 38, 58, 84]. However, they only consider homogeneous strategies and overlook the data imbalance issues.
Resources disaggregation. The idea of disaggregating storage from compute resources has been well-studied in cloud computing [74, 31, 53] and utilized in many AI applications [3, 42]. Similar concepts have also gained attention in large Transformer model inference, where disaggregating the prefilling and decoding phases allows for better throughput [67, 34, 92]. In our system, we adopt a similar design philosophy by separating the storage-heavy optimization phase from the compute-intensive propagation phase, enabling flexible execution and transitions of parallel strategies.
Intermediate representations of parallel strategies. Graph-based intermediate representations are widely utilized in deep learning systems [1, 50, 41, 32]. By analyzing and optimizing these graphs, models can discover enhanced partitioning strategies and parallel schemes. Hydraulis extends this concept by simultaneously deducing and expressing graphs for different strategies. Using subgraphs, it distinctly highlights the unique communication patterns of each strategy. Besides, any graph-based optimization techniques can be applied to Hydraulis, allowing for further refinement for each strategy.
10 Conclusion and Future Work
In this work, we addressed the often-overlooked issues of data sampling and packing imbalances in existing training systems, and introduced Hydraulis, a novel system that leverages parallel computing and data management co-designs to optimize the training performance of large Transformer models. Empirical results show that Hydraulis can achieve up to a 2.66×2.$6$ increase in throughput compared to existing systems.
As a possible future direction, we intend to expand our strategy space by integrating more parallel methods, as our system designs are fully compatible with more forms of parallelism. We also plan to tackle data imbalance challenges in diverse domains (e.g., in the video domain, the significant variation in sequence lengths has already attracted considerable attention [14, 55, 75]). We believe our work can be adapted to a broader range of scenarios and applications.
Appendix A Data imbalance
In this section, we will provide more evidence to support claims about data imbalance in §3. Firstly, we show that increasing the batch size does not really eliminate the variation in maximum sequence length across iterations. Subsequently, we conduct experiments to examine whether the quadratic time complexity of the attention mechanism impacts end-to-end performance.
Figure 14: Curve of maximum sequence length across iterations. We sample 1.6M tokens from the CommonCrawl dataset per iteration and track the longest sequence.
A.1 Larger Batch Size
In our experiment, we use 64 GPUs, processing 100K tokens per iteration. Intuitively speaking, as we scale up to 1024 GPUs, the number of tokens per iteration should increase to 1.6M. To examine whether a larger batch size would resolve the inter-iteration imbalance, we further conduct an experiment to sample 30 consecutive iterations on the CommonCrawl dataset, with 1.6M tokens per iteration. Throughout this process, we observe that the maximum sequence length still fluctuates, and falls between 15K and 32K. While this fluctuation is less pronounced compared to the scenario of using 100K tokens per iteration, it still indicates an apparent inter-iteration imbalance.
A.2 End-to-end Influence of Attention
To better demonstrate the impact of the quadratic time complexity of the attention mechanism on end-to-end training, we train the 7B model. Sequences are packed in mini-batches using the first-fit-decreasing bin packing algorithm [66], with a maximum length of 32K. Each mini-batch contains 100K tokens, and we manually ensure that they will form 4 packed sequences (if not, we just skip that iteration for this demonstration experiment). Figure 15 shows the results. Despite carefully controlling tokens and the number of packed sequences, training time still fluctuates by about 20% in both settings. This variation primarily stems from differences in the computational cost of the attention mechanism, while other computational operations and communication overhead remain relatively stable.
Appendix B Cost Model
In this section, we elaborate on the memory model $\texttt{MaxLen}(P)$, which determines the maximum sequence length, and the latency model $\texttt{T}(l,P)$, which estimates the propagation time for a specific sequence length $l$, given a parallel scheme $P$.
Figure 15: Latency breakdown and fluctuation when training LLaMA 7B on 8 Nvidia A800 GPUs. The left uses pipeline parallel (PP) without data parallel (DP), while the right uses data parallel (DP) without pipeline parallel (PP).
B.1 Memory Model
Consider a model with hidden size $H$, number of layers $\overline{\textup{layers}}$, and vocab size $V$, and there are $N$ available GPUs. During the propagation phase, we utilize the parallel scheme $TP,PP\right>$, while the optimization phase employs a fully sharded strategy. For a sequence length $l$ (i.e., micro-batch shape $[1,l]$), memory usage is calculated as follows. (a) For activations, the memory cost per micro-batch is $\frac{\overline{\textup{layers}}}{PP}\times\frac{l\times H}{TP}\times A$, where $A$ is a constant related to the model architecture, determined through profiling. The cumulative memory peak of all micro-batches (accumulated $PP$ times in the first pipeline) is $\overline{\textup{layers}}\times\frac{l\times H}{TP}\times A$. (b) For model states, given that some are fully sharded while others are related to $TP,PP\right>$, the total memory usage of all layers is $\frac{\overline{\textup{layers}}\times H^{2}}{N}\times B\times\alpha+\frac{\overline{\textup{layers}}}{PP}\times\frac{H^{2}}{TP}\times B\times\left(1-\alpha\right)$. Here, $B$ is a constant related to the model structure, obtainable via profiling, and $\alpha$ is associated with the distribution and precision of model states. For instance, if the optimization phase uses 32-bit optimizer states and parameters, while the propagation phase uses 16-bit parameters and gradients, then $\alpha=\frac{32\times 3}{32\times 3+16\times 2}=\frac{3}{4}$. Similarly, for the embedding table, the memory cost is $\frac{H\times V}{N}\times 16\times\alpha+\frac{H\times V}{TP}\times 16\times\left(1-\alpha\right)$. The total memory peak is the sum of these costs, forming a linear function of $l$. Then, the x-coordinate of this function at the y-coordinate corresponding to GPU memory capacity minus a safety margin provides the maximum sequence length $\texttt{MaxLen}(P)$ supported by $P=\left<TP,PP\right>$. It’s important to note that we can apply a similar modeling approach when incorporating other forms of parallel methods, such as context parallel and sequence parallel.
B.2 Latency Model
The latency is directly modeled as $\texttt{T}\left(l,P_{k}\right)=a_{k}\times l^{2}+b_{k}\times l+c_{k}$, where $a_{k},b_{k},c_{k}$ are coefficients related to $P_{k}$. The quadratic term reflects the quadratic scaling of attention computation with sequence length, while the linear term accounts for other computations and activation communications. The constant term represents additional overheads such as kernel launch delays. By profiling for $P_{k}$ and adjusting $l$, we fit the curve and determine these coefficients.
Appendix C Enumeration Pruning
(a) Efficiency comparison.
(b) Bubble comparison.
Figure 16: (a) LLaMA 13B training efficiency across different parallel schemes, measured in tokens per second. (b) Latency changes pre- and post-sequence packing for two equal-length sequences when $PP=2$, "F" and "B" indicate forward and backward propagation, respectively.
In this section, we will further discuss two trade-offs during packing as a supplement. Based on this, we will perform pruning optimization to avoid some unnecessary enumeration of the number of micro-batches.
  1. 1.
    Less packing, smaller micro-batch size, lower GPU utilization: As shown in Figure 16(a), we assess the GPU’s efficiency in processing sequences of specific lengths by dividing the propagation latency by the number of tokens in the sequence. Our observations indicate that efficiency approaches optimal only when the sequence length exceeds a certain threshold, which is related to intra-op parallel (e.g., TP) but independent of inter-op parallel (e.g., PP). This is due to GEMMs (General Matrix Multiplications) [64]: intra-op parallel reduces the computational load of single matrix multiplication, while inter-op parallel only reduces the number of matrix multiplications. When sequences are short or the intra-op parallel degree is large, the scale of a single matrix multiplication becomes small, leading to under-utilization of the GPU’s SMs and significant kernel launch overhead [63]. Therefore, insufficient packing on some parallel schemes may result in many under-utilized micro-batches, reducing the performance.
  2. 2.
    More packing, smaller number of micro-batches, higher pipeline bubble rate: As shown in Figure 16(b), packing inevitably reduces the number of micro-batches, increasing the bubble rate. In the example shown, there is about a 0.25×0.25$\times$0.25 × performance slowdown. Therefore, packing sequences to the memory limit or the mini-batch maximum is not necessarily optimal.
Based on these two observations, we propose that the number of micro-batches should be neither too large nor too small. Otherwise, it could lead to low GPU utilization or a high bubble rate, respectively. Therefore, we first profile each parallel scheme and introduce the value $\texttt{UtilLen}\left(P\right)$, representing the fully utilized sequence length of $P$ at which its efficiency curve reaches a certain threshold (0.85 in our experiments). This provides an upper bound for the number of micro-batches. Additionally, the maximum sequence length supported by the pipeline configured with $P$ (i.e., $\texttt{MaxLen}\left(P\right)$) offers a good lower bound. Denote the sequences dispatched to the pipeline with parallel scheme $P$ as $l_{1},\ldots,l_{U}$ and $V$ as the number of micro-batches, then the enumeration of $V$ can be pruned from the original $V\in[1,U]$ to:
$V\in\left[\max\left(\frac{\sum\limits_{i=1}^{U}l_{i}}{\texttt{MaxLen}(P)},1\right),\min\left(\frac{\sum\limits_{i=1}^{U}l_{i}}{\texttt{UtilLen}(P)},U\right)\right]$
Appendix D Details of Sequence Dispatching across Pipelines
In this section, we begin by deriving Eq. (2). We then describe the approximation algorithm for solving the sequence dispatching problem across pipelines (i.e., Eq. (3)) and compare it to the precise solution.
D.1 Derivation of the Lower Bound
Input: Sequence lengths $\{l_{1},\ldots,l_{B}\}$. Maximum strategy indices $\{J_{1},\ldots,J_{B}\}$. Parallel schemes of the pipelines $\left\{{P}_{1},\ldots,{P}_{D}\right\}$. Number of trials $T$;
Output: Optimal cost $O_{\text{best}}$. Dispatch matrix $M_{\text{best}}$;
1
2Initialize $O_{\text{best}}\leftarrow\infty$, $M_{\text{best}}\leftarrow\emptyset$;
3
4for trial $=1$ to $T$ do in parallel
5       Randomize permutation $\pi$ to shuffle sequences;
6       Initialize $C_{j},E_{j},m_{i,j}\leftarrow 0$ , $\forall i\in[1,B],j\in[1,D]$;
7
8      for $k=1$ to $B$ do
9             $i\leftarrow\pi_{k}$, $O_{\text{min}}\leftarrow\infty$, $j^{*}\leftarrow-1$;
10
11            for $j=1$ to $J_{i}$ do
12                   $l_{\text{max}}\leftarrow\max(l_{j},\max_{i}\{m_{i,j}\times l_{i}\})$;
13                   $C_{j}^{\prime}\leftarrow C_{j}+\texttt{T}\left(l_{i},{P}_{j}\right)$;
14                   $E_{j}^{\prime}\leftarrow\texttt{T}\left(l_{\text{max}},{P}_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$;
15                   $O_{\text{max}}\leftarrow\max(C_{j}^{\prime}+E_{j}^{\prime},C_{k}+E_{k}\text{ for }k\neq j)$;
16
17                  if $O_{\text{max}}<O_{\text{min}}$ then
18                         $O_{\text{min}}\leftarrow O_{\text{max}}$, $j^{*}\leftarrow j$;
19
20            $m_{i,j^{*}}\leftarrow 1$, $C_{j^{*}}\leftarrow C_{j^{*}}^{\prime}$, $E_{j^{*}}\leftarrow E_{j^{*}}^{\prime}$;
21
22      $O_{\text{trial}}\leftarrow\max_{j}\{C_{j}+E_{j}\}$;
23
24      if $O_{\text{trial}}<O_{\text{best}}$ then
25             $O_{\text{best}}\leftarrow O_{\text{trial}}$, $M_{\text{best}}\leftarrow\{m_{i,j}\}_{B\times D}$;
26
27return $O_{\text{best}}$, $M_{\text{best}}$;
Algorithm 1 Greedy-based approximate solver for sequence dispatching across pipelines.
Consider Eq. (1) for the $d$-th pipeline. Let the optimal solution be $o^{*}_{i,j}$, then the optimal objective can be split into two terms:
$A$ $=\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}\right)\times(\texttt{PP}(P_{d})-1)}$
$B$ $=\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}\right)\times V}$
For $A$, consider the longest dispatched sequence length, denoted as $l_{i_{\star}}=\max_{1\leq i\leq B}\left\{m_{i,d}\times l_{i}\right\}$. Since $\sum_{1\leq j\leq V}{o^{*}_{i_{\star},j}}=1$ and $o^{*}_{{\star},j}\in\left\{0,1\right\}$, there must exist a $j_{\star}$ such that $o^{*}_{i_{\star},j_{\star}}=1$. Thus, we have:
$A$ $\geq\sum\limits_{1\leq i\leq U}{o^{*}_{i,j_{\star}}\times\texttt{T}\left(l_{i},P_{d}\right)}\times(\texttt{PP}(P_{d})-1)$
$\geq o^{*}_{i_{\star},j_{\star}}\times\texttt{T}\left(l_{i_{\star}},P_{d}\right)\times(\texttt{PP}(P_{d})-1)$
$\geq\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{d}\right)\times(\texttt{PP}(P_{d})-1)$
For $B$, it is evident that:
$\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}$ $\leq\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}\right)},\forall j\in\left[1,V\right]$
$\implies B$ $=\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}\right)\times V}$
$\geq\sum\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o^{*}_{i,j}\times\texttt{T}\left(l_{i},P_{d}\right)}\right)}$
$=\sum\limits_{1\leq i\leq B}m_{i,d}\times\texttt{T}(l_{i},P_{d})$
By summing these two terms, the dispatching lower bound for the $d$-th pipeline is obtained:
$\small\sum_{i=1}^{B}m_{i,d}\times\texttt{T}(l_{i},P_{d})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,d}\times l_{i}\right\},P_{d}\right)\times\left(\texttt{PP}\left(P_{d}\right)-1\right)$
D.2 Approximate Solver
[…]
D.3 Performance Comparison
Figure 17: Comparison of estimated propagation latency and actual end-to-end latency between greedy dispatching and ideal dispatching. The experiments use a 32B model on the CommonCrawl dataset, with 64 GPUs. Each iteration handles 200,000 tokens with a maximum sequence length of 32K.
Strategy Proposal Ideal Dispatching Greedy Dispatching Packing
64 GPUs & 200K tokens per iteration 2.03s 40.21s 5.46s 10.15s
1024 GPUs & 3.2M tokens per iteration 107.21s >1000s 38.56s 13.46s
Table 2: Solving time for each subproblem. The time cost for ideal dispatching, greedy dispatching, and packing is listed for 10 iterations combined. Hydraulis ultimately chooses the greedy dispatching scheme over the ideal dispatching scheme due to its faster speed.
LLaMA 7B 8 GPUs 16 GPUs
F1 F2 & F3 F1 F2 & F3
DeepSpeed DP2SP4 SP8 DP4SP4 DP2SP8
Megatron-LM DP2TP2PP2 TP4PP2 DP4TP2PP2 DP2TP4PP2
HotSPa DP2TP2PP2, DP8 TP4PP2, DP4TP2PP2, DP2TP4PP2 DP4TP2PP2, DP16 DP2TP4PP2, DP4TP2PP2, DP16
LLaMA 13B 16 GPUs 32 GPUs
F1 F2 & F3 F1 F2 & F3
DeepSpeed DP4SP4 DP2SP8 DP8SP4 DP4SP8
Megatron-LM DP4TP4 DP2TP8 DP8TP4 DP4TP8
HotSPa DP4TP4, DP8TP2 DP2TP8, DP4TP4, DP8TP2 DP8TP4, DP16TP2 DP4TP8, DP8TP4, DP16TP2
LLaMA 32B 32 GPUs 64 GPUs
F1 F2 & F3 F1 F2 & F3
DeepSpeed DP8SP4 + AC DP4SP8 + AC DP16SP4 + AC DP8SP8 + AC
Megatron-LM DP2TP8PP2 DP2TP16 DP4TP8PP2 DP4TP16
HotSPa DP2TP8PP2, DP4TP4PP2 DP2TP16, DP2TP8PP2, DP4TP4PP2 DP4TP8PP2, DP8TP4PP2 DP4TP16, DP4TP8PP2, DP8TP4PP2
Table 3: Optimal training configurations. DP refers to the data parallel degree, TP to the tensor parallel degree, PP to the pipeline parallel degree, SP to the Ulysses sequence parallel degree, and AC indicates the use of activation checkpointing. For HotSPa, there are 2-3 distinct strategies.
We analyze the gap between our proposed dispatching algorithm, Alg. 1 (referred to as greedy dispatching), and the optimal solution to the original optimization problem, Eq. (3) (referred to as ideal dispatching). As illustrated in Figure 17, we compare the ratios of estimated propagation latency (i.e., the optimal objective given by the algorithm) and actual end-to-end latency by dividing the greedy dispatching results by the ideal dispatching results. The estimations closely align, and the actual end-to-end latency ratios fall between 0.9 and 1.1. This indicates that there is no significant performance difference between the two approaches. Considering the subsequent time analysis in Appendix E, we ultimately opt for greedy dispatching due to its faster solution speed and superior scalability.
Appendix E Time Cost and Scalability
In this section, we discuss the computation times of all mathematical models and their scalability. The experiments are conducted on the CommonCrawl dataset using a 32B model configuration with 64 GPUs, a maximum sequence length of 32K, and 200K tokens per iteration.
Table 2 presents the solving time for various components of our mathematical model. To enhance computational efficiency, we employ multi-processing for all components except the strategy proposal. The strategy proposal process relies on a dynamic programming algorithm, which is inherently sequential and, therefore, not suitable for parallelization. For the parallelized components, we utilize 10 CPU processes, concurrently solving the corresponding problem for 10 iterations, and record the resulting latency. To further study the scalability of our planning algorithms, apart from running real end-to-end training on the existing testbed, we also simulate a scaled-up scenario with 1024 GPUs, where the number of tokens per iteration is proportionally increased to 3.2 million.
For the strategy proposal, its computational complexity scales linearly with the number of GPUs in the cluster. Consequently, the time cost increases significantly from 64 GPUs to 1024 GPUs. However, since this component only executes once, its time cost— remaining under 2 minutes— is acceptable.
For the ideal dispatching (i.e., Eq. (3)), while the time cost is acceptable at 64 GPUs, scaling to 1024 GPUs results in superlinear growth. This is due to the complexity being proportional to both the number of pipelines and the number of sequences per iteration— both of which grow dramatically at larger scales. In contrast, greedy dispatching (i.e., Alg. 1) offers a significantly faster alternative, effectively controlling the growth in time cost. This makes it a more practical choice for large-scale deployments.
For the packing (i.e., Eq. (1)) part, we measure the time required to solve all the ILPs. Each ILP addresses the allocation of sequences within a single pipeline, and its scale remains independent of both the number of GPUs and the total number of sequences per iteration. Moreover, since these ILPs are completely independent of each other, they can be solved in parallel, resulting in a relatively stable solving time across different scales.
Overall, considering the training latency of Hydraulis in our experiments, the time cost for a single iteration is approximately 5.5 seconds, leading to a total of around 55 seconds for 10 iterations. Regardless of whether the cluster consists of 64 GPUs or 1024 GPUs, the computation time of our algorithms can be effectively overlapped with GPU execution. In the implementation of Hydraulis, we employ a dedicated process for training and adopt a producer-consumer model to coordinate planning and training. Specifically, the planning processes, acting as producers, solve the sequence dispatching and packing problems, generating execution plans. Meanwhile, the training process, acting as the consumer, retrieves and executes the plans for the current iteration. This design ensures that GPUs remain fully saturated with training workloads throughout execution, thereby achieving efficient utilization of computational resources.
Appendix F Optimal Training Configurations
Table 3 illustrates the manually tuned configurations that achieve the best training performance under different scenarios in § 8.2.
  • For scenario F1, with a maximum sequence length of 16K, two strategies are employed: (a) The first strategy handles sequences of 4K to 16K. (b) The second strategy handles sequences less than 4K.
  • For scenarios F2 and F3, with a maximum sequence length of 32K, three strategies are employed: (a) The first strategy handles sequences of 16K to 32K. (b) The second strategy handles sequences of 4K to 16K. (c) The third strategy handles sequences less than 4K.

### v2

@@
-Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences
-Abstract
-To optimize large Transformer model training, efficient parallel computing and advanced data management are essential.
-However, current methods often assume a stable and uniform training workload, neglecting imbalances in data sampling and packing that can impede performance.
-Specifically, data sampling imbalance arises from uneven sequence length distribution of the training data, while data packing imbalance stems from the discrepancy between the linear memory complexity and quadratic time complexity of the attention mechanism.
-To address these imbalance issues, we develop Hydraulis, which jointly optimizes the parallel strategies and data assignment.
-For one thing, we introduce large model training with dynamic heterogeneous parallel strategies in response to the sequence length variations within and across training iterations.
-For another, we devise a two-stage data assignment approach, which strikes a good balance in terms of the training workloads both within and across model replicas.
-Empirical results demonstrate that Hydraulis outperforms existing systems by 1.32-2.66$\times$.
-1 Introduction
-In recent years, large Transformer models [80] have become pivotal in fields such as natural language processing [21, 9, 70], computer vision [10, 23, 56], video processing [7, 2, 26], and speech recognition [22, 4, 30].
-Their success relies on pre-training models with vast parameters on extensive datasets [43, 25].
-However, this process is both time-consuming and costly, making the acceleration of large model training a pressing concern for both academia and industry.
-To move towards Artificial General Intelligence (AGI), input data are condensed into sequences of tokens to achieve a unified representation [60, 23], and models are growing larger in size [43].
-Thus, large model training takes as input variable-length sequences and has substantial memory consumption.
+Hydraulis: Balancing Large Transformer Model Training via Co-designing Parallel Strategies and Data Assignment
+Abstract.
+[…]
+1. Introduction
+In recent years, large Transformer models (Vaswani et al., 2017) have become pivotal in fields such as natural language processing (Devlin et al., 2019; Brown et al., 2020; Raffel et al., 2020), computer vision (Carion et al., 2020; Dosovitskiy et al., 2021; Liu et al., 2021), video processing (Bertasius et al., 2021; Arnab et al., 2021; Fan et al., 2021), speech recognition (Dong et al., 2018; Baevski et al., 2020; Gulati et al., 2020) and intelligent agents (Zhou et al., 2024; Zhang et al., 2024b; Xu et al., 2025; Xi et al., 2025).
+Their success relies on pre-training models with vast parameters on extensive datasets (Kaplan et al., 2020; Dubey et al., 2024).
+Figure 1. Sequence length distribution of two popular open-sourced datasets. We display the number of sequences and tokens within each length range, with both datasets exhibiting high variance.
+Within this landscape, two key challenges arise.
+(a) To achieve better model ability, efforts have been made to scale models to larger sizes (Kaplan et al., 2020), making large model training substantially time-consuming and memory-intensive.
+(b) Meanwhile, to move towards Artificial General Intelligence (AGI), input data are often condensed into sequences of tokens to achieve a unified representation (Mikolov et al., 2012; Dosovitskiy et al., 2021; Du et al., 2024).
+However, this representation naturally exhibits length variability (de Vries, 2023).
+For instance, Figure 1 presents the sequence length distribution after tokenizing two popular large language models (LLMs) training datasets—CommonCrawl and GitHub—revealing that the sequence lengths are significantly diverse in the real world.
+Thus, large model training takes as input variable-length data.
 To address this, approaches from two perspectives have been developed.
-(a) From a parallel computing perspective, methods like data parallel [52, 87, 88, 5], tensor parallel [77, 62, 46], pipeline parallel [35, 44, 61, 49, 69, 33], sequence and context parallel [36, 54, 51, 8, 29, 27] have been proposed, each with trade-offs in latency and memory usage, and are often combined for better performance.
-(b) From a data management perspective, data packing techniques are widely used in the training of both language [47, 66, 6, 83, 48] and vision [20, 15] models to minimize redundant computations by combining sequences for simultaneous processing.
+(a) From a parallel computing perspective, methods like data parallel (Li et al., 2020; You et al., 2020; Zhang et al., 2024a; Bai, 2022), tensor parallel (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022), pipeline parallel (Huang et al., 2019; Kim et al., 2020; Narayanan et al., 2021a; Lamy-Poirier, 2023; Qi et al., 2023; Harlap et al., 2018; Guan et al., 2024), sequence and context parallel (Jacobs et al., 2023; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023; Gu et al., 2024; Fang and Zhao, 2024; Wang et al., 2025a) have been proposed, and are often combined to train large-scale models.
+Data parallel operates at the granularity of training pipelines (i.e., data parallel groups containing model replicas that are capable of independent forward and backward propagation) and thus can typically scale by adding more pipelines, while other parallel methods (i.e., tensor, pipeline, sequence and context parallel) are optimized within a single training pipeline to balance pipeline latency and memory usage.
+(b) From a data management perspective, data packing techniques—combining shorter sequences into longer ones for simultaneous processing—are widely used in the training of both language (Krell et al., 2022; NVIDIA, 2023d; Bai et al., 2024; Wang et al., 2024b; Kundu et al., 2024) and vision (Dehghani et al., 2023; Choudhury et al., 2024) models to minimize redundant computations.
 Despite these advancements, existing training systems typically leverage these two perspectives in a decoupled manner.
-(a) In parallel computing, most systems overlook the sequence length variation and adopt a static, homogeneous parallel strategy [77, 72, 91, 38], which deploys all training pipelines with the same parallel scheme throughout the training.
-[^1] In this work, we use the term “training pipeline” (or “pipeline”) to denote a model replica that can carry out forward and backward propagation individually (i.e., a data parallel group).
-We use “parallel scheme” to denote the parallel configuration of a pipeline (each pipeline does not necessarily employ any parallel methods, i.e., it can be deployed on only one GPU).
+(a) In parallel computing, most systems overlook the sequence length variation and adopt a static, homogeneous parallel strategy (Shoeybi et al., 2019; Rasley et al., 2020; Zheng et al., 2022; Jia et al., 2019), which deploys all training pipelines (a.k.a.
+model replicas) with the same parallel scheme throughout the training (e.g., all pipelines have identical tensor parallel and pipeline parallel degree).
+[^1] In this work, we use “parallel scheme” to denote the parallel configuration of a pipeline (each pipeline does not necessarily employ any parallel methods, i.e., it can be deployed on only one GPU).
 Given the available GPUs, a “parallel strategy” describes the collection of parallel schemes of all pipelines deployed.
 (b) In data management, existing works usually specify a maximum sequence length (a.k.a.
-context length), pack the original sequences into multiple packed ones, and distribute the packed sequences evenly across machines [66, 48].
-In short, current works enforce a static, homogeneous parallel strategy and a fixed-length packing on variable-length data.
-Nevertheless, as investigated in §3, such a decoupled design fails to take account of two key imbalances.
-Firstly, sequences in the real world vary in length by nature, following a long-tail distribution in general [19].
+context length), pack the original sequences into multiple packed ones that do not exceed the maximum length (termed max-length packing), and distribute the packed sequences evenly across machines (NVIDIA, 2023d; Kundu et al., 2024).
+In short, current works enforce a static, homogeneous parallel strategy and a max-length packing on variable-length data.
+However, our empirical analysis (details shown in §3) reveals that this decoupled design overlooks two critical imbalances.
+Firstly, sequences in the real world substantially vary in length by nature (de Vries, 2023).
 This leads to the data sampling imbalance, manifesting as variable sequence lengths within iterations and inconsistent maximum lengths across iterations.
-When employing a homogeneous parallel strategy, we must ensure the parallel scheme is memory-saving enough (e.g., with a high tensor parallel degree) to support the maximum sequence length, which compromises the efficiency of short sequences.
+When employing a homogeneous parallel strategy, we must ensure the parallel scheme is memory-saving enough (e.g., with a high context parallel or tensor parallel degree) to support the maximum sequence length, which compromises the efficiency of short sequences.
 The second is the data packing imbalance.
 Due to the attention mechanism in Transformer, there is a discrepancy between the quadratic time complexity and the linear memory complexity w.r.t.
-the sequence length [80, 18, 66].
-However, existing works adopt fixed-length packing, achieving memory balance but leaving the training workloads imbalanced both within each pipeline and among different pipelines.
-Motivated by this, this work manages to address these imbalance issues by jointly optimizing the parallel and packing strategies, putting forward co-designs in both the parallel computing and data management perspectives.
+the sequence length (Vaswani et al., 2017; Dao et al., 2022; NVIDIA, 2023d).
+However, existing works adopt max-length packing, achieving memory balance but leaving the training workloads imbalanced within and across pipelines.
+Motivated by this, we manage to address these imbalance issues by jointly optimizing the parallel and packing strategies, putting forward co-designs in both the parallel computing and data management perspectives.
 To this end, we introduce Hydraulis, an innovative system designed for efficient training of large Transformer models.
@@
 (II) Disaggregating optimization and propagation with subgraphs.
-We present a novel disaggregation of the optimization and propagation phases in distributed training, which disentangles the divergent model partitioning brought by dynamic heterogeneous strategies.
-We further introduce the subgraph abstraction to unify the representation of sophisticated communication needs during propagation (for activations) and between optimization and propagation (for parameters and gradients).
+To support training with dynamic heterogeneous parallel strategies, we present a novel disaggregation of the optimization phase (i.e., synchronizing gradients and updating the model) and propagation phase (i.e., executing forward and backward propagation in a pipelined manner) in distributed training.
+This approach maintains consistent model states sharding for optimization, and can transit to arbitrary heterogeneous parallel strategy for propagation.
+We further introduce subgraphs (i.e., subsets of the complete computation graph) to unify the representation of sophisticated communication needs during propagation (for activations) and optimization-propagation interactions (for parameters and gradients).
 This allows for flexible combinations of parallel schemes and seamless transitions between parallel strategies.
 (III) Two-stage sequence assignment.
-Given a mini-batch of sequences and an arbitrary strategy, we devise a two-stage management of sequences to eliminate the data packing imbalance.
+Given a mini-batch of sequences and an arbitrary strategy, we devise a two-stage approach to eliminate the data packing imbalance.
 In particular, we formulate an optimization problem that aims to determine how packed sequences can be balanced within each pipeline.
@@
 (IV) Data distribution-aware strategy proposal.
-Guided by the sequence assignment and for practical considerations, we further propose preparing the candidate strategies that are promising for the given training task in advance, in order to accelerate the strategy enumeration and selection for each iteration.
+Guided by the sequence assignment and for practical considerations, we further propose preparing the candidate strategies that are promising in advance, in order to accelerate the strategy enumeration and selection for each iteration.
 It takes the overall sequence length distribution of the entire dataset into account, and proposes the candidate strategies via a dynamic programming process.
-(V) Evaluation. Extensive experiments demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over state-of-the-art training systems, while also offering better scalability.
-2 Preliminaries
-2.1 Data Sampling and Data Packing
+[…]
+2. Preliminaries
+2.1. Data Sampling and Data Packing
 During the training of a Transformer model, each iteration samples a mini-batch of sequences with a pre-defined threshold on the total number of tokens, performs forward and backward propagation on the mini-batch, and updates the model.
-Besides, the model is typically trained with a pre-defined context length, representing the maximum number of tokens it can process in a single input [68].
+Besides, the model is typically trained with a pre-defined context length, representing the maximum number of tokens it can process in a single input (Pawar et al., 2024).
 Any sequence exceeding this limit is typically truncated to fit within the context length.
@@
 This variability introduces challenges when constructing batches, as sequences need to be aligned to a common length for efficient batch processing.
-Traditionally, this issue is addressed through data padding, where shorter sequences are padded with special tokens (e.g., [PAD] tokens) to match the length of the longest one [21, 80].
-Although padding ensures uniformity in sequence length and is straightforward to implement, it leads to unnecessary computation on the padded tokens, results in wasted computational resources.
-Figure 1: Left: Illustration of padding and packing.
+Traditionally, this issue is addressed through data padding, where shorter sequences are padded with special tokens (e.g., [PAD] tokens) to match the length of the longest one (Devlin et al., 2019; Vaswani et al., 2017).
+Although padding ensures uniformity in sequence length and is straightforward to implement, it leads to unnecessary computation on the padded tokens, resulting in wasted computational resources.
+Figure 2.
+Left: Illustration of padding and packing.
 Right: Comparison of the attention latency during LLaMA2 7B training on Nvidia A800 GPUs with tensor parallel (TP).
@@
 (b) Disaggregate optimization and propagation.
-(c) Subgraphs instantiation.
-Figure 2: "O, P, G" represents optimizer states, parameters and gradients, respectively.
-"AG, RS, S, R" represents all-gather, reduce-scatter, send and receive, respectively.
-"To" represents the data type transfer operator (from 32- to 16-bit floating points), and "LNK" represents the link operator designed by us.
+Figure 3.
+“O, P, G” represents optimizer states, parameters and gradients, respectively.
+“AG, RS, S, R” represents all-gather, reduce-scatter, send and receive, respectively.
+“To” represents the data type transfer operator (from 32- to 16-bit floating points), and “LNK” represents the link operator designed by us.
 (a) The overall workflow of the model states sharding technique combined with mixed-precision training.
 (b) Our proposed optimization-propagation disaggregation technique, which decouples the distributed strategies for the optimization phase and propagation phase, allowing for arbitrary heterogeneous parallel strategies during propagation.
-(c) The instantiation of Pull Parameter and Communicate Activation subgraphs during forward propagation.
-This involves inserting custom operators (e.g., send-receive, all-gather, LNK) to support arbitrary propagation strategies and optimization-propagation interactions.
-As shown in Figure 1, an alternative to padding is data packing [47], which seeks to maximize computational efficiency by grouping sequences of varying lengths into a single longer sequence, thereby eliminating the need for padded tokens.
-To prevent cross-contamination between the packed sequences, positional embeddings and the attention mask must be adjusted to preserve the integrity of the attention mechanism [6, 83, 48].
-Fortunately, modern techniques such as FlashAttention v2 have integrated support for block diagonal attention, which natively handles variable-length sequences (i.e., varlen) [18, 17].
-Moreover, the varlen interface ensures that the attention complexity scales as $\sum{S_{i}}^{2}$, rather than $(\sum{S_{i}})^{2}$, where $S_{i}$ represents the length of the $i$-th original sequence being packed [66].
-This can significantly reduce the computational overhead when using data packing, making it a more practical and efficient approach.
-2.2 Parallel Methods, Schemes, and Strategies
+As shown in Figure 2, an alternative to padding is packing (Krell et al., 2022), which seeks to maximize computational efficiency by grouping sequences of varying lengths into a single longer sequence, thereby eliminating the need for padded tokens.
+To prevent cross-contamination between the packed sequences, positional embeddings and the attention mask must be adjusted to preserve the integrity of the attention mechanism (Bai et al., 2024; Wang et al., 2024b; Kundu et al., 2024).
+Fortunately, modern techniques such as FlashAttention v2 have integrated support for block diagonal attention, which natively handles variable-length sequences (i.e., varlen) (Dao et al., 2022; Dao, 2023).
+Moreover, the varlen interface ensures that the attention complexity scales as $\sum{S_{i}}^{2}$, rather than $(\sum{S_{i}})^{2}$, where $S_{i}$ represents the length of the $i$-th original sequence being packed (NVIDIA, 2023d).
+This significantly reduces the computational overhead of data packing, making it a more practical and efficient approach.
+2.2. Parallel Methods, Schemes, and Strategies
 As models grow in both parameter size and computational complexity, distributed training across multiple devices becomes essential.
 Below we introduce popular parallel methods for large Transformer model training.
-Data parallel.
-In data parallel (DP), the model is replicated across multiple devices, with each device processing a different subset of the input data [52, 87, 88, 5].
+Data parallel
+In data parallel (DP), the model is replicated across multiple devices, with each device processing a different subset of the input data (Li et al., 2020; You et al., 2020; Zhang et al., 2024a; Bai, 2022).
 After computing the gradients locally, devices must synchronize their model updates, typically through a gradient reduction process, to ensure consistency across devices.
-Parallel scheme of one pipeline.
+Parallel scheme of one pipeline
 In this work, we call each data parallel group one training pipeline, and the parallel scheme describes how a pipeline is deployed.
@@
 Data from a mini-batch is further split into micro-batches, which flow through the pipeline in a staged manner.
-(When training Transformer models, each micro-batch accounts for one packed sequence.) Techniques such as micro-batch scheduling and gradient accumulation are often employed, allowing multiple micro-batches to be processed efficiently [35, 44, 61, 49, 69, 33].
+(When training Transformer models, each micro-batch accounts for one packed sequence.) Techniques such as micro-batch scheduling and gradient accumulation are often employed, allowing multiple micro-batches to be processed efficiently (Huang et al., 2019; Kim et al., 2020; Narayanan et al., 2021a; Lamy-Poirier, 2023; Qi et al., 2023; Harlap et al., 2018; Guan et al., 2024).
 Unlike inter-op parallel, there are several intra-op parallel methods for a pipeline.
-Tensor parallel (TP) partitions each layer across multiple GPUs and exchanges intermediate results (e.g., via all-gather and reduce-scatter operations) to complete the forward and backward passes [77, 62, 46].
-Additionally, sequence parallel and context parallel have been explored to handle extremely long sequences, where a single sequence is partitioned across multiple devices to avoid out-of-memory (OOM) errors [36, 54, 51, 8, 29, 27].
-Parallel strategy.
+Tensor parallel (TP) partitions each layer across multiple GPUs and exchanges intermediate results (e.g., via all-gather and reduce-scatter operations) to complete the forward and backward passes (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022).
+Additionally, sequence parallel (SP) and context parallel (CP) have been explored to handle extremely long sequences, where a single sequence is partitioned across multiple devices to avoid out-of-memory (OOM) errors, and all-to-all or ring-style peer-to-peer communication would be necessary to perform the attention mechanism over the partitioned sequences (Jacobs et al., 2023; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023; Gu et al., 2024; Fang and Zhao, 2024).
+Parallel strategy
 Building on the parallel scheme of a single pipeline, the parallel strategy refers to the collection of parallel schemes of all pipelines.
 However, existing works mainly leverage the same parallel schemes for all pipelines, resulting in a homogeneous parallel strategy.
-Such a homogeneous design fails to take account of the imbalances caused by data sampling and data packing, as discussed in §3.
-2.3 Model States Sharding
+Such a homogeneous design fails to take account of the imbalances caused by data sampling and data packing, which will be discussed in §3.
+2.3. Model States Sharding
 To reduce memory consumption during training, model states sharding techniques (a.k.a.
-ZeRO [71] and FSDP [89, 88]) are widely used.
+ZeRO (Rajbhandari et al., 2020) and FSDP (Zhao et al., 2023; Zhang et al., 2024a)) are widely used.
 These methods extend standard data parallel (DP) by partitioning model states (i.e., parameters, gradients, and optimizer states) across multiple devices, rather than replicating them across all data parallel groups (i.e., pipelines).
-Figure 3: The dot plot on the left shows the fluctuation of maximum sequence length, emphasizing inter-iteration imbalance, while the bar chart on the right presents the distribution of sequence lengths within an iteration (20th iteration), highlighting intra-iteration imbalance.
-The sharding technique is combined with mixed-precision training [59, 65], which uses both 16-bit floating point (FP16 or BF16) and 32-bit floating point (FP32) precisions to balance efficiency and numerical stability.
+The sharding technique is combined with mixed-precision training (Micikevicius et al., 2018; NVIDIA, 2023c), which uses both 16-bit floating point (FP16 or BF16) and 32-bit floating point (FP32) precisions to balance efficiency and numerical stability.
 Figure 3(a) illustrates an example with each pipeline deployed with a tensor parallel degree of 2.
 The 32-bit optimizer states and parameters (which consume more memory but are critical for ensuring convergence) are sharded across GPUs, while the 16-bit parameters and gradients (which are more memory-efficient and computationally faster) are maintained during propagation.
-As aforementioned, current works mainly adopt a homogeneous parallel strategy, so they usually align the sharding with the strategy, allowing distributed state transformations to be easily achieved via all-gather and reduce-scatter operations.
-However, since such transformations can overlap with computation, it is also feasible to adopt more nuanced, non-aligned parameter pulling and gradient pushing using customized communication strategies, which will be elaborated in §5.2.
-3 Imbalances Analysis
+As aforementioned, current works mainly adopt a homogeneous parallel strategy, so they usually align the sharding with the parallel strategy, allowing distributed state transformations to be easily achieved via all-gather and reduce-scatter operations.
+However, since such transformations can overlap with computation, it is also feasible to adopt more nuanced, non-aligned parameter pulling and gradient pushing using customized communication strategies, which is depicted in Figure 3(b) and will be elaborated in §5.2.
+2.4. Graph-based Representations
+The execution of a deep learning model can be represented as a computation graph—a directed acyclic graph (DAG) where nodes correspond to operators (e.g., matrix multiplication, attention) defined by users and edges indicate data dependencies.
+Modern deep learning frameworks, such as PyTorch (Li et al., 2020) and TensorFlow (Abadi et al., 2016), adopt this graph-based representation to encapsulate execution logic.
+Following this convention, we employ the same representation but extend it with a subgraph abstraction to enable dynamic heterogeneous strategies, which will be presented in §5.3.
+2.5. Limitations of State-of-the-art Systems
+State-of-the-art (SOTA) training systems, including Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022) and DeepSpeed (Rasley et al., 2020; Jacobs et al., 2023), have incorporated diverse parallel methods (e.g, Megatron-LM supports DP, TP, PP and CP, while DeepSpeed offers DP and SP) and model states sharding techniques (e.g., Megatron-LM implements ZeRO-1, which shards the optimizer states and the FP32 parameters, while DeepSpeed supports ZeRO 1-3 to allow further sharding FP16/BF16 gradients and parameters).
+However, they are limited to using one static parallel strategy throughout the training task.
+A more recent work, HotSPa (Ge et al., 2024), enables the use of multiple pre-defined parallel strategies and supports switching between them within an iteration.
+This enables data in different sequence lengths within an iteration to utilize different parallel strategies while accumulating gradients for a global update.
+Nonetheless, as will be analyzed in §3, various aspects of data imbalance necessitate support for dynamic heterogeneous strategies.
+However, all of these systems lack this capability, as they are constrained to homogeneous parallel strategies (i.e., all pipelines use the same parallel scheme) and require model states sharding to align with the parallel strategy.
+This is due to the following reasons: (a) Existing designs do not support parameter retrieval and gradient aggregation across heterogeneous pipelines, as they rely on homogeneous communication primitives (i.e., all-gather and reduce-scatter) but lack heterogeneous ones.
+(b) Balancing workload across heterogeneous pipelines is challenging due to the absence of fine-grained data assignment mechanisms and strategy design methods that take data distribution into account.
+To address (a), Hydraulis adopts an optimization-propagation disaggregation architecture to facilitate sophisticated parameter retrieval and gradient aggregation, and introduces subgraphs to support arbitrary heterogeneous strategies (§5).
+To address (b), it proposes a two-stage sequence assignment method for load balancing (§6), and pre-generates distribution-aware heterogeneous parallel strategies before training (§7).
+Together, these innovations overcome the limitations of current training systems and improve training efficiency under data imbalance conditions (§8).
+3. Imbalances Analysis
 In this section, we analyze the imbalance issues encountered during data sampling and packing in training on real-world datasets, and discuss potential optimization opportunities.
-To ease the analysis, we exemplify the imbalance issues by training the LLaMA 7B model with the CommonCrawl [78] and GitHub [16] datasets over 8 GPUs.
-The context length is 32K and the batch size is set as 100K for each iteration.
-3.1 Data Sampling Imbalance
+To ease the analysis, we exemplify the imbalance issues by training the LLaMA 13B model with the CommonCrawl (TIIUAE, [n.
+d.]) and GitHub (CodeParrot, [n.
+d.]) datasets over 16 GPUs, with 32K context length and 100K tokens per iteration.
+Figure 4.
+The dot plot on the left shows the fluctuation of maximum sequence length, emphasizing inter-iteration imbalance, while the bar chart on the right presents the distribution of sequence lengths within an iteration (20th iteration), highlighting intra-iteration imbalance.
+3.1. Data Sampling Imbalance
 We first concentrate on the imbalance in terms of data sampling.
@@
 In the following, we analyze the two issues, respectively.
-Intra-iteration imbalance.
+Intra-iteration imbalance
 The uneven sequence length distribution within an iteration can be attributed to the inherent property of data.
 In particular, sequences in the real world generally follow a long-tail distribution, where short sequences dominate and long sequences are rare.
-Existing studies have also observed such a distribution in many popular training datasets [19].
+Existing studies have also observed such a distribution in many popular training datasets (de Vries, 2023).
 This imbalance poses both challenges and opportunities when it comes to optimizing overall performance.
 The challenges stem from the trade-offs between memory usage and throughput offered by different numbers of GPUs and parallel schemes.
-As shown in Figure 4, increasing the number of GPUs generally leads to a lower throughput per GPU but allows for a longer maximum supportable sequence length.
-Meanwhile, under the same number of GPUs, different parallel schemes (e.g., PP, TP and their combinations) also give divergent trade-offs.
-As aforementioned in §2.2, existing works utilize all GPUs in the cluster uniformly with a homogeneous parallel strategy tailored for the longest context length, compromising the training efficiency for memory.
-Nevertheless, opportunities arise when adapting the trade-offs to sequences of varying lengths: short sequences, which are more frequent, can be processed using faster but more memory-intensive parallel schemes, while the fewer long sequences can be assigned to slower yet more memory-efficient parallel schemes.
+As shown in Figure 5, increasing the number of GPUs generally leads to a lower throughput per GPU but allows for a longer maximum sequence length.
+Meanwhile, under the same number of GPUs, different parallel schemes (e.g., PP, TP and their combinations) also have divergent trade-offs (e.g., replacing PP with TP generally supports longer sequences but at the cost of reduced per-GPU throughput).
+As aforementioned in §2.2, existing works utilize all GPUs in the cluster uniformly with a homogeneous parallel strategy tailored for the longest context length, compromising the overall training efficiency for memory.
+Nevertheless, combined with the length imbalance within a mini-batch, such trade-offs become particularly invaluable: short sequences can be processed using faster but more memory-intensive parallel schemes, while the long sequences can be assigned to slower yet more memory-efficient parallel schemes.
 Therefore, adopting heterogeneous parallel strategies— where non-unique parallel schemes are applied on different proportions of GPUs based on different sequence lengths— enables better performance (§5).
-Figure 4: Memory and throughput trade-offs.
-"OOM" indicates out-of-memory, and "N/A" indicates not available.
+Figure 5.
+Memory and throughput trade-offs.
+“OOM” indicates out-of-memory, and “N/A” indicates not available.
 The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs.
-Throughput is measured in tokens per second.
-Inter-iteration imbalance.
+Throughput is measured in tokens per second and maximum sequence lengths are given by $\texttt{MaxLen}(\cdot)$ (see §6.1 and Appendix C.1 (hyd, 2025)).
+Inter-iteration imbalance
 While intra-iteration imbalance arises from the distribution of sequence lengths within a single mini-batch, another challenge comes from the distribution variations across mini-batches.
-Particularly, due to the long-tail distribution, rare long sequences may not appear in every iteration (even with a large batch size, the maximum sequence length still varies, as detailed in Appendix A.1).
+Particularly, due to the long-tail distribution, rare long sequences may not appear in every iteration (even with a large batch size on a 1024-GPU cluster, the maximum sequence length still varies, as detailed in Appendix A.1 (hyd, 2025)).
 This leads to fluctuations in data distribution and maximum sequence lengths across iterations.
@@
 To address this, we propose to dynamically adjust the parallel strategy based on the characteristics of each mini-batch to better align the workload, thereby enhancing overall performance across iterations (§5).
-Figure 5: The numbers of original sequences (classified by their lengths) in different micro-batches after packing (each micro-batch accounts for one packed sequence).
-3.2 Data Packing Imbalance
+Discussions
+[…]
+Additionally, a prior work, HotSPa (Ge et al., 2024), addresses the data sampling imbalance in a different way—it performs length-based training within each iteration instead of over the entire dataset, in order to preserve the model convergence.
+By dynamically switching between expert-tuned, homogeneous parallel strategies for different sequence lengths, HotSPa accumulates gradients and performs a single global update per iteration.
+However, it lacks support for heterogeneous strategies and finer-grained data assignment, and frequent strategy switching introduces additional overhead.
+We will empirically compare HotSPa and our work in §8.
+Figure 6.
+Training and evaluation loss of different sampling methods.
+We subsample 250M tokens for training and 5M for validation from the GitHub dataset, and train a 13B model on 16 GPUs with 32K context length and 100K tokens per iteration.
+3.2. Data Packing Imbalance
 We then focus on the imbalance issues of data packing.
-As shown in Figure 5, we observe that due to varying sequence lengths, there are significant differences in the distributions of original sequences within the packed sequences, both among different micro-batches within the same pipeline and across different pipelines.
-In the following, we will analyze the impact of each aspect individually.
-Intra-pipeline imbalance.
-Since the memory consumption of each micro-batch is linear to the total length after packing, existing works commonly adopt fixed-length packing methods, like bin packing [86, 66], to balance memory usage.
-However, due to the quadratic time complexity of attention [80, 18, 66], the computation latency can hardly be balanced with fixed-length packing— for example with Figure 1 again, attention latency can be nearly 20×20$\times$20 × smaller when packing 32 sequences of 1K compared to a single 32K sequence.
-In pipeline parallel (PP), this leads to uneven execution times across micro-batches, yet common pipeline scheduling, such as 1F1B [33] or zero-bubble [69], requires uniform execution time of micro-batches to minimize pipeline bubbles.
-This highlights the importance of taking account of both the linear memory complexity and quadratic time complexity, rather than considering solely the memory as in fixed-length packing.
+As illustrated in Figure 7, we observe significant distribution differences in the original sequences within packed sequences (i.e., micro-batches) due to varying sequence lengths.
+These variations occur both within micro-batches of the same pipeline and across different pipelines, resulting in execution time variance.
+Below, we analyze each aspect’s impact separately.
+Intra-pipeline imbalance
+Since the memory consumption of each micro-batch is linear to the total length after packing, existing works commonly adopt max-length packing methods, like bin packing (Yao, 1980; NVIDIA, 2023d), to balance memory usage.
+However, due to the quadratic time complexity of attention (Vaswani et al., 2017; Dao et al., 2022; NVIDIA, 2023d), the computation latency can hardly be balanced in this way— for example with Figure 2 again, attention latency can be nearly $20\times$ smaller when packing 32 sequences of 1K compared to a single 32K sequence.
+In pipeline parallel (PP), this leads to uneven execution times across micro-batches, yet common pipeline scheduling, such as 1F1B (Harlap et al., 2018) or zero-bubble (Qi et al., 2023), requires uniform execution time of micro-batches to minimize pipeline bubbles.
+This highlights the importance of taking account of both the linear memory complexity and quadratic time complexity, rather than considering solely the memory as in max-length packing.
 To address this, we devise a nuanced packing strategy that considers the variation in original sequence lengths, striking a balance in terms of the execution time across different micro-batches within the memory constraint (§6.1).
-Inter-pipeline imbalance.
-Although different pipelines execute concurrently, the training latency is dictated by the slowest pipeline due to the need for gradients synchronization.
-Simply packing sequences to a fixed length has a similar issue of unbalanced computation load.
+Inter-pipeline imbalance
+Although different pipelines execute concurrently, the training latency is dictated by the slowest pipeline due to the need for gradient synchronization.
+Simply packing sequences to a fixed maximum length has a similar issue of unbalanced computation load.
 In some cases, if the sequence number after packing cannot be evenly distributed across pipelines, a more severe imbalance will arise.
@@
 This innovative dispatching paradigm ensures a more balanced load across pipelines, before organizing sequences into micro-batches (§6.2).
-4 Overview
+Figure 7.
+The numbers of original sequences (classified by their lengths) in different micro-batches after packing (each micro-batch accounts for one packed sequence) and their corresponding time cost.
+The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs.
+4. Overview
 […]
@@
 To tackle the data packing imbalance, given a mini-batch of sequences and a strategy, we devise a two-stage assignment (§6).
-The first stage determines how to dispatch these sequences of varying lengths to different pipelines to solve the inter-pipeline imbalance.
+The first stage determines how to dispatch sequences of varying lengths to different pipelines to solve the inter-pipeline imbalance.
 Then, the second stage meticulously packs the sequences within each pipeline to form different micro-batches, addressing the intra-pipeline imbalance.
@@
 […]
-5 System Designs
-In this section, we start by mathematically defining our heterogeneous parallel strategies, then introduce the system techniques proposed to support these strategies and seamless transitions between them.
-5.1 Dynamic Heterogeneous Strategies
+5. System Designs
+In this section, we start by mathematically defining our heterogeneous parallel strategies, then introduce the system techniques proposed to support these strategies and the dynamic transitions between them.
+5.1. Dynamic Heterogeneous Strategy Design
+Definitions
 […]
+$\mathbb{P}=\big\{P|$ $\texttt{TP}(P)\in\mathbb{D}_{T},\texttt{PP}(P)\in\mathbb{D}_{P},\texttt{CP}(P)\in\mathbb{D}_{C},$
+$\texttt{TP}(P)\times\texttt{PP}(P)\times\texttt{CP}(P)\leq N\big\},$
+where $\mathbb{D}_{T}=\mathbb{D}_{C}=\{1,2,\cdots,N\}$, $\mathbb{D}_{P}=\left\{v\in\mathbb{N}^{+}|v\text{ divides }\overline{\textup{layers}}\right\}$, $N$ is the number of GPUs, and $\overline{\textup{layers}}$ is the number of layers.
 Next, we define our heterogeneous strategy space as the linear span of the parallel scheme space: $\mathbb{S}=\text{span}(\mathbb{P})$, meaning that each heterogeneous strategy can be represented as a linear combination of parallel schemes: $S=\sum_{i=1}^{K}d_{i}\times P_{i}$, which indicates that $S$ consists of $\sum_{i=1}^{K}d_{i}$ pipelines of $K$ different parallel schemes, with $d_{i}$ pipelines deployed with $P_{i}$.
-Strategy execution and transition.
+Figure 8 shows some example parallel schemes ($P$) and heterogeneous strategies ($S$) when the number of GPUs ($N$) is 8.
+Strategy execution and transition
 During each training iteration, given a mini-batch of sequences, we select the most suitable $S$ to execute.
 [^3] The selection in each iteration is based on the two-stage sequence assignment (§6), and we propose the candidate strategies by analyzing the sequence length distribution of the entire dataset (§7).
-The pipelines will be assigned with non-overlapping sequences and execute the forward and backward propagation concurrently, which is termed the propagation phase.
-Then, akin to traditional data parallel (DP), all pipelines synchronize the model gradients and perform updates, which is termed the optimization phase.
-Since different iterations inevitably select non-unique strategies, strategy transitions between iterations are necessary.
-System challenges.
-However, enabling the system to support various pipelines, each configured with a different $P$, is complex due to the need to manage diverse communication patterns across devices.
-Furthermore, dynamically transitioning between strategies at the iteration level presents additional difficulties, as the model states mappings on devices can vary significantly across different $S$.
-To address these issues, we develop an optimization-propagation disaggregation approach (§5.2) and a subgraph-based abstraction (§5.3).
-These system designs successfully meet our requirements of training with dynamic, heterogeneous strategies.
-5.2 Optimization-propagation Disaggregation
-As illustrated in Figure 2(b), unlike traditional methods that employ a uniform distributed training strategy, we separate the optimization phase from the propagation phase, decoupling the GPU’s role as a distributed storage from its computational role during propagation.
-Specifically, this approach maintains a constant partitioning for FP32 parameters and optimizer states, while allowing any parallel strategy to be used during propagation, enabling the system to support dynamic heterogeneous strategies and seamless transitions between them.
-[^3] We use a fully sharded strategy in optimization phase, yet any other zero-redundancy sharding strategies are also feasible.
-When propagation requires access to FP32 parameters and optimizer states, we perform pull and push interactions, which correspond to the retrieval of parameters and the synchronization and scattering of gradients, respectively.
-These interactions occur only at the beginning and end of a training iteration, referred to the steps 1-2 and steps 4-5 of Figure 2(b).
-To effectively manage pull and push interactions, we develop a specialized LNK (link) operator to handle complex interaction patterns.
-As depicted on the right side of Figure 7, this operator utilizes three core primitives: send, receive, and local movement, centered around a fundamental unit called the mutual slice, which represents the common partitioning granularity when model shardings interact between the propagation and optimization phases.
-Given that the optimization states are partitioned on each GPU without redundancy, the pull and push schemes can be uniquely determined using these three primitives on the non-duplicate mutual slices.
-In particular, assume there are $N$ GPUs, the optimization phase employs a fully sharded strategy, and the propagation phase uses a strategy $S=\sum_{i=1}^{K}d_{i}\times P_{i}$ with $D=\sum_{i=1}^{K}d_{i}$ pipelines.
+The pipelines will be assigned different sequences and execute forward and backward passes concurrently to compute gradients from the corresponding parameters.
+This process is termed the propagation phase (Figure 3(a)3(b), step 3).
+Then, similar to traditional data parallelism (DP), all pipelines synchronize the model gradients and perform updates at the end of each iteration to obtain the new parameters and optimizer states.
+This process is called the optimization phase (Figure 3(a)3(b), step 6).
+Since different iterations inevitably select non-unique strategies, strategy transitions between iterations are necessary (e.g., in Figure 8, iteration 8 uses strategy $S5$, while iteration 9 transitions to strategy $S2$).
+Challenges
+To support the execution and transition of arbitrary heterogeneous strategies, the conventional approach to handling model states must be revised.
+As shown in Figure 3(a), for current systems that employ a static, homogeneous parallel strategy, it is common practice to shard the model parameters and optimizer states according to the employed homogeneous strategy, and leverage all-gather and reduce-scatter operations to retrieve the parameters and synchronize the gradients, respectively.
+However, this approach falls short in supporting dynamic heterogeneous strategies, as model partitioning varies significantly across pipelines.
+For example, in Figure 3(b), pipelines with different TP degrees produce parameters and gradients with divergent partitioning granularities (i.e., tensors are split into non-uniform slices across GPUs).
+Consequently, it is infeasible to discover a single communication pattern to accomplish the parameter retrieval or gradient synchronization.
+This limitation drives our development of the LNK (link) operator, which supports parameter retrieval and gradient synchronization in an optimization-propagation disaggregated manner (§5.2), and we further introduce subgraphs to encapsulate diverse communication patterns (§5.3).
+Furthermore, since gradient synchronization requires all pipelines to participate and is constrained by the slowest one, pipeline heterogeneity also poses significant load-balancing challenges, which we will address in §6.
+5.2. Optimization-Propagation Disaggregation
+As illustrated in Figure 3(b), we separate the optimization phase from the propagation phase, decoupling the GPU’s role as a distributed storage from its computational role during propagation.
+Specifically, we keep the partitioning of FP32 parameters and optimizer states fixed for the optimization phase [^4] We use a fully sharded strategy in the optimization phase, yet any other zero-redundancy sharding strategies are also feasible., while allowing an arbitrary parallel strategy to be used in the propagation phase.
+In such a disaggregated design, to ensure correct parameter partitioning during propagation as well as correct gradient partitioning for optimizer update, we establish two types of interaction: (a) pull, which retrieves the parameters from the optimization phase to an arbitrary heterogeneous parallel strategy (Figure 3(b), steps 1–2); (b) push, which synchronizes and scatters gradients back to the optimization phase (Figure 3(b), steps 4-5).
+Compared to the all-gather and reduce-scatter in the conventional design that requires identical, homogeneous parallel strategy in the two phases (Figure 3(a), steps 1-2 and 4-5), our pull and push abstractions enable the use of arbitrary heterogeneous parallel strategies in the propagation phase as well as the seamless transition between different strategies.
+Figure 9.
+Diagram of RS (reduce-scatter) and LNK (link) operations during push.
+Pink and blue denote slices in the propagation and optimization phases, respectively, while green represents the mutual slice (the finest-grained slice shared by both).
+The partitioning granularity is 4 in the propagation phase and 6 in the optimization phase (after reduce-scatter).
+Thus, the mutual slice has a granularity of 12 (i.e., the least common multiple of 4 and 6).
+To effectively manage pull and push interactions, we develop a specialized LNK (link) operator to handle complex communication patterns.
+As depicted on the right side of Figure 9, this operator utilizes three core primitives: send, receive, and local movement, all operating on a fundamental unit called the mutual slice, which represents the finest-grained data partition.
+These primitives enable dynamic resharding of parameters and gradients between propagation and optimization phases.
+For instance, in Figure 9, GPU1 generates a gradient slice during propagation, which is further partitioned into two finest-grained mutual slices (7 and 8).
+Since GPU0 requires mutual slice 7 during the optimization phase for model updates, it triggers a send to transfer mutual slice 7 from GPU1 to GPU0.
+Because optimization states are uniquely partitioned across GPUs (without redundancy), the use of these primitives is deterministic.
+Mathematically, assume there are $N$ GPUs, the optimization phase employs a fully sharded strategy, and the propagation phase uses a strategy $S=\sum_{i=1}^{K}d_{i}\times P_{i}$.
 Let $\overline{TP}_{\text{max}}=\max_{i\in[1,K]}\texttt{TP}(P_{i})$.
@@
 Similarly, for the push interaction, as shown on the left side of Figure 9, a reduce-scatter is first performed at the corresponding location to synchronize the gradient.
-Then the partitioning granularity of the mutual slice is $\mathtt{LCM}\left(N,D\times\overline{TP}_{\text{max}}\right)$, with the additional $D$ accounting for the gradient’s reduce-scatter.
+Then the partitioning granularity of the mutual slice is $\mathtt{LCM}\left(N,D\times\overline{TP}_{\text{max}}\right)$, where the additional factor $D=\sum_{i=1}^{K}d_{i}\times\texttt{CP}(P_{i})$ accounts for gradients reduce-scatter across all model replicas.
 The LNK can be established still using the send, receive and local movement primitives.
-Figure 7: Diagram of RS (reduce-scatter) and LNK (link) operations during push.
-Pink indicates the gradient, blue the optimizer states, and green the mutual slice.
-The optimizer states are partitioned with a granularity of 4, and the gradient after reduce-scatter has a granularity of 6.
-Thus, the mutual slice has a granularity of 12 (i.e., the least common multiple of 4 and 6).
-Compared to training with homogeneous strategies, introducing the LNK operator incurs additional overhead to resolve the discrepancy between strategies.
+Notably, the LNK mechanism operates independently of $\texttt{PP}(P_{i})$, as PP doesn’t partition individual parameters into slices.
+From the perspective of heterogeneous strategy execution, the LNK operator resides in different stages across training pipelines.
+Consequently, communication occurs only after the slowest pipeline reaches the LNK operator (the balancing across pipelines will be addressed in §6).
+It is noteworthy that our optimization-propagation disaggregation is fundamentally a generalization of current model states sharding techniques (i.e., ZeRO (Rajbhandari et al., 2020) and FSDP (Zhao et al., 2023))— when using homogeneous strategies, the LNK operator in the pull will be translated into an all-gather operation, and the LNK operator in the push will be pruned (leaving a reduce-scatter operation), which are the same as current techniques.
+For heterogeneous strategies, introducing the LNK operator incurs additional overhead to resolve the discrepancy between strategies.
 However, the send and receive communication can be overlapped with computation during propagation, and the local movement incurs only minimal latency, so the overhead is small (as evaluated in §8.4).
-Last but not least, our design offers great generalization capabilities— when the optimization and propagation use the same strategy, the LNK operator in the pull will be translated into an all-gather operation, and the LNK operator in the push will be pruned (leaving a reduce-scatter operation).
-5.3 Subgraphs
-To tackle the challenge of representing diverse heterogeneous strategies and their complex optimization-propagation interactions, we introduce subgraphs— subsets of the complete computation graph.
+Last but not least, theoretically speaking, the communication during pull and push operations is bounded by the model size and does not scale with the number of GPUs.
+And the worst-case communication volume is exactly the same as that of current model states sharding techniques.
+Furthermore, computation-communication overlap can be achieved when each micro-batch in the pipeline with $P_{i}$ contains more than $\texttt{TP}(P_{i})\times\texttt{CP}(P_{i})\times\frac{F}{B}$ tokens, where $F$ is the GPU’s half-precision compute capability (FLOPS), and $B$ is the slowest interconnect bandwidth (bytes/sec).
+This condition is independent of the model size or the number of GPUs.
+Due to the space constraint, detailed proof and analysis are provided in Appendix B (hyd, 2025).
+5.3. Subgraphs
+Although optimization-propagation disaggregation offers the ability to resolve the communication patterns needed for each heterogeneous strategy, it still poses the challenge of how to manage such diverse and complex optimization-propagation interactions, given that there would be numerous candidate heterogeneous strategies.
+To tackle this challenge, we introduce subgraphs, which are subsets of the complete computation graph.
 The key observation is that different strategies only have unique patterns in some specific operators, while most user-defined computational operators remain unchanged.
 Therefore, subgraphs are developed to encapsulate these unique operators from different strategies for efficient representation, management and scheduling.
-Figure 8: Scheduling subgraphs with pipeline parallel (PP).
-In the first PP stage, Pull Parameter subgraphs are scheduled immediately after the last operators requiring the parameter, and Push Gradient subgraphs follow right after the operators that produce or accumulate the gradient.
-In subsequent PP stages, Pull Parameter and Push Gradient are grouped at the beginning and end, respectively, to utilize inherent PP bubbles.
-This scheduling scheme effectively hides communication latency within the Pull Parameter and Push Gradient subgraphs.
+Figure 10.
+The instantiation of Pull Parameter and Communicate Activation subgraphs during forward propagation.
+This involves inserting custom operators (e.g., send-receive, all-gather, LNK) to support arbitrary propagation strategies and optimization-propagation interactions.
 In particular, we define three subgraph types: Pull Parameter, Push Gradient, and Communicate Activation.
 Operators for interactions between optimization and propagation phases, like LNK, are placed in the first two, while activation-related communication operators (e.g., all-gather and reduce-scatter for TP, send-receive for PP) are included in the third.
-Subgraphs instantiation.
-As illustrated at the bottom of Figure 2(c), users start by defining various computational operators to describe the training model.
+Subgraphs instantiation
+As illustrated in Figure 10, users start by defining various computational operators to describe the training model.
 We then identify all potential points in the computational graph where different heterogeneous strategies might intervene and preserve several empty subgraphs (e.g., between transformer modules for potential pipeline parallel, before and after matrix multiplication for potential tensor parallel).
@@
 We then determine how to instantiate all predefined subgraphs with appropriate operators.
-For example, as depicted at the top of Figure 2(c), a data transfer operator and a strategy-specific LNK operator are added to instantiate a Pull Parameter subgraph.
+For example, as depicted at the top of Figure 10, a data transfer operator and a strategy-specific LNK operator are added to instantiate a Pull Parameter subgraph.
 This instantiation process ultimately forms a complete graph for a specific $S$, enabling different GPUs to extract and execute their respective operators.
-Subgraphs scheduling.
+Subgraphs scheduling
 During execution, we optimize overlap by controlling the execution order of subgraphs and computational operators.
@@
 As illustrated in Figure 11, by using a typical cudaStream to execute the subgraphs and interleaving them with computational operators, we enable communication within subgraphs to overlap with computation, thereby reducing latency.
-6 Two-stage Sequence Assignment
+6. Two-stage Sequence Assignment
 In this section, we present our two-stage sequence assignment approach.
@@
 In each iteration, we apply the assignment to all candidate strategies, estimate the propagation latency, and select the optimal one.
-6.1 Sequence Packing within Pipeline
+Figure 11.
+Scheduling subgraphs with pipeline parallel (PP).
+In the first $P$ stage, Pull Parameter subgraphs are scheduled immediately after the last operators requiring the parameter, and Push Gradient subgraphs follow right after the operators that produce or accumulate the gradient.
+In subsequent $P$ stages, Pull Parameter and Push Gradient are grouped at the beginning and end, respectively, to utilize inherent $P$ bubbles.
+This scheduling scheme effectively hides communication latency within the Pull Parameter and Push Gradient subgraphs.
+6.1. Sequence Packing within Pipeline
 We first introduce the packing strategy within each pipeline.
@@
 To pack these sequences into multiple micro-batches, there are two essential terms to determine: the number of micro-batches to craft (denoted as $V$) and how each sequence should be assigned.
-Cost Model.
-For a given $P$, there is a maximum sequence length it can support, denoted as $\texttt{MaxLen}(P)$ (i.e., sequence exceeding this length will cause OOM on $P$), which can be determined in advance using our memory model (detailed in Appendix B.1).
-Additionally, we introduce a latency model, $\texttt{T}(l,P)$, which estimates the time taken to run a sequence of length $l$ on $P$ for a single forward and backward propagation of a single pipeline stage (detailed in Appendix B.2).
-Solving for one $V$.
+Cost Model
+Following prior works on automatic parallel training (Zheng et al., 2022; Li et al., 2025b; Lin et al., 2025; Zhao et al., 2025), we develop two profiling-based cost models to estimate memory usage and latency.
+For a given $P$, there is a maximum sequence length it can support, denoted as $\texttt{MaxLen}(P)$ (i.e., sequence exceeding this length will cause OOM on $P$), which can be determined in advance using our memory model (detailed in Appendix C.1 (hyd, 2025)).
+Additionally, we introduce a latency model, $\texttt{T}(l,P)$, which estimates the time taken to run a sequence of length $l$ on $P$ for a single forward and backward propagation of a single pipeline stage (detailed in Appendix C.2 (hyd, 2025)).
+Similar to prior works, re-profiling is required for different model architectures but is performed as a one-shot process per architecture.
+This overhead is worthwhile given the significant speedup in large-scale training.
+Moreover, this process is independent of the cluster size, as scaling the cluster only increases the number of pipelines, without affecting the set of valid $P$ or the sequence length $l$ used during training.
+Solving for one $V$
 The problem is easy to solve if the number of micro-batches $V$ is given.
-To be specific, as mentioned in §2.1, by employing the varlen functionality of FlashAttention v2 [18, 17], the latency of a packed sequence can be approximated as the sum of the latency of the individual original sequences.
+Specifically, as mentioned in §2.1, by employing the varlen functionality of FlashAttention v2 (Dao et al., 2022; Dao, 2023), the latency of a packed sequence can be approximated as the sum of latency for each original sequence.
 Meanwhile, to address the intra-pipeline imbalance, our goal is to minimize the duration of the longest-running micro-batch.
 Thus, we can formulate the following problem on top of the minimum sum partition (MSP) problem:
-$\operatorname*{arg\,min}_{o_{i,j}}$ $\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o_{i,j}\times\texttt{T}\left(l_{i},P\right)}\right)\times(\texttt{PP}(P)-1+V)}$ (1)
+(1) $\operatorname*{arg\,min}_{o_{i,j}}$ $\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o_{i,j}\times\texttt{T}\left(l_{i},P\right)}\right)\times(\texttt{PP}(P)-1+V)}$
 s.t. $\sum\limits_{1\leq j\leq V}{o_{i,j}}=1,o_{i,j}\in\left\{0,1\right\},\forall i\in\left[1,U\right]$
@@
 Compared to the MSP problem, the problem defined in Eq.
-(1) additionally ensures that the length of each packed sequence must not exceed $\texttt{MaxLen}(P)$, and we multiply the term $\texttt{PP}(P)-1+V$ to represent the propagation latency of the entire pipeline.
-Enumeration of $V$.
-To locate the optimal number of micro-batches $V^{*}$, we can simply enumerate the value for $V\in[1,U]$, solve Eq.
-(1), and choose the best one.
+(1) additionally ensures that the length after packing must not exceed $\texttt{MaxLen}(P)$, and the term $\texttt{PP}(P)-1+V$ represents the propagation latency of the entire pipeline.
+Enumeration of $V$
+To locate the optimal number of micro-batches $V^{*}$, we can simply enumerate the value for $V\in[1,U]$, solve Eq. (1), and choose the best one.
 Note that Eq.
 (1) is an Integer Linear Programming (ILP) problem, but its complexity is only related to the number of sequences assigned to $P$ (i.e., $U$) and the number of micro-batches (i.e., $V$).
-These do not increase with cluster size, and the enumeration can be done in parallel, ensuring fast problem-solving even on large-scale clusters (detailed in Appendix E).
-Additionally, we can prune unnecessary enumerations, where a too-small $V$ results in a high bubble rate and a too-large $V$ leads to excessive kernel launch overheads.
-Detailed pruning methods are provided in Appendix C.
-6.2 Sequence Dispatching across Pipelines
+These do not increase with cluster size, and the enumeration can be done in parallel, ensuring fast problem-solving even on large-scale clusters (detailed in Appendix F (hyd, 2025)).
+6.2. Sequence Dispatching across Pipelines
 Next, we consider how the sequences in a mini-batch should be dispatched across the pipelines to address the inter-pipeline imbalance.
+A naïve approach is to bucket sequences by length.
+However, given the long-tail distribution of sequence lengths (where long sequences are rare, as shown in Figure 1), pipelines configured for long sequences may receive too few assignments during iterations dominated by short sequences.
+To improve load balancing, we need a more flexible approach that allows short sequences to be processed by any pipeline.
 Consider a specific strategy $S=P_{1}+\ldots+P_{D}$ satisfying $\texttt{MaxLen}\left(P_{1}\right)\geq\ldots\geq\texttt{MaxLen}\left(P_{D}\right)$.
@@
 Fortunately, it is feasible to estimate the lower bound of latency.
-In particular, let $m_{i,j}\in\{0,1\}$ indicate whether the $i$-th sequence is assigned to $P_{j}$, then the lower bound of running time of the $j$-th pipeline can be expressed as (detailed derivation in Appendix D.1):
-$\small\sum_{i=1}^{B}m_{i,j}\times\texttt{T}(l_{i},P_{j})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$ (2)
+In particular, let $m_{i,j}\in\{0,1\}$ indicate whether the $i$-th sequence is assigned to $P_{j}$, then the lower bound of running time of the $j$-th pipeline can be expressed as (detailed derivation in Appendix E.1 (hyd, 2025)):
+(2) $\small\sum_{i=1}^{B}m_{i,j}\times\texttt{T}(l_{i},P_{j})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$
 Denote Eq. (2) as $\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$, then the dispatching can be formulated as a (0,1)-ILP problem:
-$\operatorname*{arg\,min}_{m_{i,j}}$ $\max\limits_{1\leq j\leq D}\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$ (3)
+(3) $\operatorname*{arg\,min}_{m_{i,j}}$ $\max\limits_{1\leq j\leq D}\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$
 s.t. $\sum\limits_{1\leq j\leq J_{i}}{m_{i,j}}=1,\forall i\in\left[1,B\right]$
@@
 In practice, as the cluster scale expands, both will increase proportionally, making the assignment problem-solving cost unacceptable in large-scale scenarios.
-Fortunately, there are extensive studies about how to approximately solve assignment problems in polynomial time with a guaranteed error bound [24, 76, 12].
-Motivated as such, we implement an approximate solver to Eq.
-(3), which utilizes multiple random trials to greedily dispatch sequences to pipelines and ultimately select the best trial.
-Empirical results demonstrate that the approximate solver achieves nearly the same objectives as directly solving Eq.
-(3), but with greater speed.
-More details about our approximate solver can be found in Appendix D and E.
-7 Data Distribution-aware Strategy Proposal
+Fortunately, there are extensive studies about how to approximately solve assignment problems in polynomial time with a guaranteed error bound (Duan and Pettie, 2014; Shmoys and Tardos, 1993; Chen et al., 2020).
+Thus, we follow previous works to solve Eq.
+(3) approximately, which involves running multiple random trials to greedily dispatch sequences and selecting the best trial.
+Empirical results demonstrate that the approximation errors are consistently lower than 10% compared to directly solving Eq.
+(3).
+More details about our approximate solution can be found in Appendix E and Appendix F (hyd, 2025)
+7. Data Distribution-aware Strategy Proposal
 Ideally, for each iteration, we could explore all possible strategies, estimate their propagation latency via the two-stage sequence assignment, and find the optimal one.
@@
 Thus, we analyze the dataset in advance and propose a refined subset $\hat{\mathbb{S}}\subset\mathbb{S}$, which only contains the most promising candidates.
-Figure 9: Dynamic programming with non-integer values.
-Different colors represent various $P$.
-The segments indicate sequences within each dataset interval, with their lengths corresponding to latency.
-Rounding the number of $P$ requires scaling the resulting latency.
-Problem definition.
-To address both intra- and inter-iteration imbalances of data sampling, the proposal is based on two considerations.
-Firstly, since intra-iteration imbalance typically reflects the expected distribution of the dataset, we aim to preserve strategies that demonstrate an overall advantage over the entire dataset.
-Secondly, due to inter-iteration imbalance, characterized by variations in maximum sequence lengths, we aim to identify non-unique strategies across different maximum lengths.
+Problem definition
+Our proposal is based on two considerations regarding intra- and inter-iteration imbalances: First, since intra-iteration imbalance typically reflects the expected distribution of the dataset, we aim to preserve strategies that demonstrate an overall advantage over the entire dataset.
+Second, due to inter-iteration imbalance, characterized by variations in maximum sequence lengths, we aim to identify non-unique strategies across different maximum lengths.
 Then, we have the following problem.
@@
 Later we lift the restriction by continuous relaxation, enhancing the solution.
-Dynamic programming.
+Dynamic programming
 We define $t[n][l]$ as the minimum achievable propagation latency using up to $n$ GPUs to process all sequences in the dataset with lengths not exceeding $l$.
@@
 Therefore, we have the following state transition expression:
-$t[n][l]=\min\left(t[n-1][l],\min_{(k,d)\in{Cond}_{n,l},l^{\prime}<l}\left\{{T}_{k,d,l^{\prime}}\right\}\right),\ \textup{where}$ (4)
+(4) $t[n][l]=\min\left(t[n-1][l],\min_{(k,d)\in{Cond}_{n,l},l^{\prime}<l}\left\{{T}_{k,d,l^{\prime}}\right\}\right),\ \textup{where}$
 ${T}_{k,d,l^{\prime}}=\max\left(t[n-d\times\texttt{N}\left(P_{k}\right)][l-l^{\prime}],\frac{1}{d}\sum_{x\in\mathbb{D}_{(l-l^{\prime},l]}}\texttt{T}(x,P_{k})\right)$
@@
 We can then select all unique $S[N][L]$ from $L=1$ to $L=L_{max}$ to form $\hat{\mathbb{S}}$.
-Continuous relaxation.
+Figure 12.
+Dynamic programming with non-integer values.
+Different colors represent various $P$.
+The segments indicate sequences within each dataset interval, with their lengths corresponding to latency.
+Rounding the number of $P$ requires scaling the results.
+Continuous relaxation
 Based on the dynamic programming, we lift the restriction that each pipeline can only handle a sequence over a single length interval.
@@
 Consequently, the final solution $S[N][L]$ might involve having non-integer numbers, as exemplified in Figure 12.
-In such cases, we round and adjust to include all nearby integer solutions in $\hat{\mathbb{S}}$, while ensuring that the total number of GPUs does not exceed $N$ [^4] The whole algorithm’s worst-case time complexity is $O(KN^{2}L_{\text{max}}^{2})$.
-In our experiments, we set the step size of $n$ and $d$ to 0.1 and $l$ to 128, with the actual time cost detailed in Appendix E.
+In such cases, we round and adjust to include all nearby integer solutions in $\hat{\mathbb{S}}$, while ensuring that the total number of GPUs does not exceed NN.
+[^5] The whole algorithm’s worst-case time complexity is $O(KN^{2}L_{\text{max}}^{2})$.
+In our experiments, we set the step size of $n$ and $d$ to 0.1 and $l$ to 128, with the actual time cost detailed in Appendix F (hyd, 2025).
 Undoubtedly, even though the original non-integer solution achieves optimal results, rounding compromises the optimality.
@@
 Due to these limitations, we can only suggest potential strategy candidates, whereas the two-stage sequence assignment is still necessary for selecting optimal strategies.
-8 Evaluation
-Figure 10: End-to-end evaluation.
+Figure 13.
+End-to-end evaluation.
 We test the workloads of three configurations, F1-F3, in each scenario.
-"Max Seq" represents the number of tokens in the longest sequence, while "Iter" denotes the total number of tokens sampled in a single iteration.
+“Max Seq” represents the number of tokens in the longest sequence, and “Iter” denotes the number of tokens sampled (a.k.a.
+batch size) in a single iteration.
 Error bars indicate the standard deviation across iterations.
-8.1 Experimental Setup
-Environment.
-Our experiments are conducted with 8 GPU nodes, each containing 8 Nvidia A800 GPUs (64 GPUs in total).
+8. Evaluation
+Figure 14. Comparison of scalability across various factors, using the fastest one as the reference for relative changes. Shaded areas indicate the standard deviation across iterations.
+(a) Case study w.r.t parallel strategies.
+(b) Case study w.r.t sequence assignment.
+Figure 15.
+(a) Latency over 100 training iterations, with colors indicating the strategy selected for each iteration.
+(b) Comparison of load balancing between Hydraulis and Megatron-LM.
+We present the maximum and minimum propagation latency among all ranks for 5 iterations.
+8.1. Experimental Setup
+Environment
+Our experiments are conducted with 8 nodes, each containing 8 Nvidia A800 GPUs, with a total number of 64 GPUs.
 GPUs within a node are connected via NVLink with a bandwidth of 400 GB/s, and nodes are interconnected using Infini-Band with a bandwidth of 200 GB/s.
-Baselines.
-We compare Hydraulis with Megatron-LM [77, 62, 46], DeepSpeed [72, 36], and HotSPa [28].
+Baselines
+We compare Hydraulis with Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022), DeepSpeed (Rasley et al., 2020; Jacobs et al., 2023), and HotSPa (Ge et al., 2024).
 Both Megatron-LM and DeepSpeed employ a single, static homogeneous strategy throughout the training process.
@@
 However, due to the overhead of switching, HotSPa is limited to setting 3-4 strategies.
-We tune the strategies for all competitors to achieve the best performance.
-Models and datasets.
+We enumerate all competitors’ parallel strategies and select the optimal one to maximize baseline performance (see Appendix G (hyd, 2025) for details).
+Models and datasets
 Most existing large models are based on the decoder-only Transformer architecture.
-We select the widely used LLaMA2 series models and conduct experiments with 7B, 13B and 32B model sizes [79].
-For datasets, we use CommonCrawl [78] and GitHub [16], which are widely used open-source datasets and have served as the major training corpora for many prestigious Transformer models.
-Worloads and metrics.
-By default, we sample 100K tokens per iteration [^5] We consider 100K tokens per iteration for our 64-GPU cluster, scaled proportionally from the 16M used in LLaMA with an 8192-GPU cluster [25].
-with a context length of 32K.
-And the Adam optimizer [45] is used.
+We select the widely used LLaMA2 series models and conduct experiments with 7B, 13B and 32B model sizes (Touvron et al., 2023), using CommonCrawl (TIIUAE, [n.
+d.]) and GitHub (CodeParrot, [n.
+d.]), which are widely used open-source datasets and have served as the major training corpora for many prestigious Transformer models.
+Worloads and metrics
+By default, we sample 100K tokens per iteration [^6] We consider 100K tokens per iteration for our 64-GPU cluster, scaled proportionally from the 16M used in LLaMA with an 8192-GPU cluster (Dubey et al., 2024).
+with a 32K context length, and use the Adam optimizer (Kingma and Ba, 2015).
 We conduct 20 warm-up iterations and then summarize the throughput (tokens per second) or per-iteration latency over 100 iterations.
-Figure 11: Comparison of scalability across various factors, using the fastest one as the reference for relative changes. Shaded areas indicate the standard deviation across iterations.
-(a) Case study w.r.t parallel strategies.
-(b) Case study w.r.t sequence assignment.
-Figure 12: (a) Latency over 100 training iterations, with colors indicating the strategy selected for each iteration.
-(b) Comparison of load balancing between Hydraulis and Megatron-LM.
-We present the maximum and minimum propagation latency among all ranks for 5 iterations.
-8.2 End-to-end Experiments
-We first evaluate Hydraulis against three baselines across various configurations (baselines’ strategies are detailed in Appendix F).
-The results in Figure 10 demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over the best baselines.
+8.2. End-to-end Experiments
+[…]
 As DeepSpeed and Megatron-LM are limited to a single fixed strategy, they must adopt less efficient strategies to accommodate the memory consumption of long sequences, degrading the overall efficiency given most sequences are relatively short.
 Although HotSPa is capable of using multiple homogeneous strategies within an iteration to train sequences of varying lengths, it requires multiple times of model hot switching.
-This incurs significant overheads, especially when per-iteration time is short (e.g., in the “GitHub 7B 8GPUs F1” scenario, the average iteration time is merely 6.7s, while nearly 2s is spent on hot switching).
+This incurs apparent overheads, especially when per-iteration time is short (e.g., in the “GitHub 7B 8GPUs F1” scenario, the average iteration time is merely 6.7s, while nearly 2s is spent on hot switching).
+[^7] A more detailed analysis of HotSPa’s switching overhead can be found in Appendix H (hyd, 2025).
 On the contrary, Hydraulis handles variable-length sequences using dynamic heterogeneous strategies, while eliminating the overhead of strategy transition via disaggregating optimization and propagation.
-[…]
-8.3 Scalability
+Moreover, none of the baselines consider fine-grained sequence management.
+This problem is particularly severe for HotSPa— in some cases, there are only 1-2 long sequences per iteration, yet HotSPa is forced to adopt an individual strategy with all GPUs to process such long sequences, leaving a few GPUs with no sequences to process (e.g., in the “CommonCrawl 32B 64GPUs F2” scenario, HotSPa adopts $\left<TP,PP,CP\right>\times DP=\left<16,1,1\right>\times 4$ to process sequences ranging 16K-32K in length, often resulting in 3/4 or 1/2 GPUs being completely unused).
+Instead, Hydraulis, with the two-stage sequence assignment, meticulously organizes the training workloads of different sequences for better performance, as we will evaluate more in-depth later in §8.4.
+8.3. Scalability
 We then assess the scalability of different systems by examining three factors: (a) number of GPUs, (b) maximum sequence length, and (c) tokens per iteration.
-Using the LLaMA 7B model and the CommonCrawl dataset, we start with a default setup of 16 GPUs, a maximum sequence length of 32K, and 200K tokens per iteration.
-By systematically altering one factor while keeping the others constant, we gain insights into each factor’s impact.
-The results are presented in Figure 11.
-Number of GPUs.
+Using the LLaMA 7B model and the CommonCrawl dataset, we start with a default setup of 16 GPUs, 32K maximum sequence length, and 200K tokens per iteration.
+By systematically altering one factor while maintaining others, we gain insights into each factor’s impact.
+Results are presented in Figure 14.
+Additionally, we analyze how (a) data imbalance and (b) problem-solving cost scale with increasing training workloads, particularly when both the cluster size (number of GPUs) and batch size (tokens per iteration) grow proportionally.
+Assessment: number of GPUs
 As the number of GPUs increases, all systems experience a sub-linear reduction in latency due to the communication overhead introduced by parallelism.
@@
 Consequently, its latency consistently outperforms other baselines.
-Maximum sequence length.
+Assessment: maximum sequence length
 As the maximum sequence length grows, DeepSpeed and Megatron-LM resort to less efficient strategies, resulting in a significant increase in latency.
@@
 In contrast, by customizing parallel strategies for sequences of different lengths, HotSPa and Hydraulis are less sensitive to maximum sequence length changes, with Hydraulis demonstrating superior performance over HotSPa.
-Tokens per iteration.
+Assessment: tokens per iteration
 As the number of tokens per iteration increases, all systems experience an almost linear increase in latency.
 However, HotSPa and Hydraulis exhibit a slightly lower growth rate, as they take account of the long-tail distribution of sequence lengths within iterations.
-8.4 Performance Interpretation
-Figure 13: Ablation study. Error bars indicate the standard deviation across iterations. Total throughput changes as different components of Hydraulis are incrementally integrated.
+Analysis: data imbalance
+When scaled to more GPUs, each iteration typically samples more training data, which may mitigate inter-iteration imbalance because longer sequences become more likely to be selected.
+To assess this, we conduct a 1024-GPU simulation with 16M tokens sampled per iteration (detailed in Appendix A.1 (hyd, 2025)).
+The results show that the maximum sequence length fluctuates between 16K-32K, indicating that inter-iteration imbalance still remains.
+Meanwhile, intra-iteration imbalance persists due to the dataset’s inherent sequence length distribution, while intra- and inter-pipeline imbalances originate from the discrepancy between the attention mechanism’s quadratic time complexity and linear memory complexity relative to sequence length.
+Since these imbalances are scale-independent, they persist at larger scales.
+Analysis: problem-solving cost
+We also conduct simulation experiments to evaluate the scalability of problem solving when scaled to 1024 GPUs (detailed in Appendix F (hyd, 2025)), which show that the problem-solving cost is well-bounded.
+To be specific, the strategy proposal (§7) incurs acceptable overhead (¡ 2 minutes) as it executes only once prior to training.
+Meanwhile, both sequence packing (§6.1) and dispatching (§6.2) operate on CPUs and can be fully overlapped with GPU training (i.e., the problem-solving for subsequent iterations occurs concurrently with current iteration training).
+Furthermore, larger cluster sizes provide additional CPU resources to increase the concurrency of problem solving, allowing us to further amortize the solving time per iteration.
+Figure 16. Ablation study. Error bars indicate the standard deviation across iterations. Total throughput changes as different components of Hydraulis are incrementally integrated.
+8.4. Performance Interpretation
 To further evaluate the impact of the proposed techniques, we conduct experiments by training the 32B model with the CommonCrawl dataset over 64 GPUs.
-Case study.
+Case study
 Figure 15(a) shows the optimal strategies selected by Hydraulis, along with the corresponding latencies, and summarizes the frequency of different strategies over 100 iterations.
 It can be seen that: (a) the actual propagation latency closely matches our estimation in most cases, and (b) variations across iterations lead to different strategy choices.
+To understand why using different strategy choices contributes, we compare the estimated latencies of two strategies, $S_{0}$ and $S_{3}$, over different iterations.
+While $S_{3}$ achieves $1.6\times$ speedup than $S_{0}$ in most cases due to its heterogeneous design, its maximum supportable sequence length is shorter than $S_{3}$, making it inapplicable for certain iterations.
+Besides, when the sampled batch contains fewer short sequences, it’s hard for $S_{3}$ to strike a balance (i.e., many sequences can only be processed by the pipeline with a TP degree of 8 while the other pipelines receive very few sequences), making it less efficient than $S_{0}$.
+These findings highlight the advantages of heterogeneous strategies, as well as the trade-off between latency and maximum sequence length across different strategies, demonstrating the benefits of supporting dynamic heterogeneous strategies for overall optimal performance.
 […]
-Additionally, we examine the load balancing.
-Figure 12(b) shows that Megatron-LM, with fixed-length packing and even dispatching of packed sequences, results in a latency difference of 0.36-0.65$\times$ between the slowest and fastest ranks.
-In contrast, Hydraulis reduces this gap to just 0.02-0.10$\times$, verifying the significance of the two-stage sequence assignment.
-Ablation study.
+Ablation study
 Figure 16 presents an ablation study by progressively enabling different techniques in Hydraulis to evaluate their contributions: (a) Starting with a static, homogeneous parallel strategy, we dispatch sequences of varying lengths according to Eq.
@@
 (d) Finally, we enable dynamic heterogeneous strategies, utilizing all Hydraulis designs.
-The results show consistent improvements across (a)-(d), demonstrating that each component of the system plays a critical role in enhancing overall performance.
-Latency breakdown.
+The results demonstrate that, on the one hand, when no optimizations are applied (i.e., Hydraulis-Base), our performance is on par with Megatron-LM, confirming the fairness of the comparison.
+On the other hand, consistent improvements are observed across (a)-(d), highlighting the critical role each system component plays in boosting overall performance.
+Table 1.
+Breakdown of latency per iteration for different training strategies.
+“P” denotes propagation (Figure 3(b) step 3), “O” refers to optimization (Figure 3(b) step 6), and “O-P” represents the remaining overhead from the interaction between the two phases (Figure 3(b) steps 1-2 and 4-5).
+“E2E” stands for end-to-end training latency.
+Homo Hetero w/o Overlap Hetero w/ Overlap
+P 6.78s 4.29s ($1.58\times$) 4.29s ($1.58\times$)
+O 0.02s 0.02s ($1.00\times$) 0.02s ($1.00\times$)
+O-P 1.11s 1.71s ($0.69\times$) 1.18s ($0.94\times$)
+E2E 7.91s 6.02s ($1.31\times$) 5.49s ($1.44\times$)
+Latency breakdown
 We further investigate the effectiveness of optimization-propagation disaggregation compared to conventional model states sharding, as well as the benefits brought by subgraphs scheduling.
@@
 Besides, we experiment both with and without subgraphs scheduling to measure the latency improvements from overlapping pull and push interactions with computation.
-Homo Hetero w/o Overlap Hetero w/ Overlap
-P 6.78s 4.29s (1.58×1.58$\times$1.58 ×) 4.29s (1.58×1.58$\times$1.58 ×)
-O 0.02s 0.02s (1.00×1.$0$) 0.02s (1.00×1.$0$)
-O-P 1.11s 1.71s (0.69×0.69$\times$0.69 ×) 1.18s (0.94×0.94$\times$0.94 ×)
-E2E 7.91s 6.02s (1.31×1.31$\times$1.31 ×) 5.49s (1.44×1.$4$)
-Table 1: Breakdown of latency per iteration for different training strategies.
-"P" denotes propagation (Figure 2(b) step 3), "O" refers to optimization (Figure 2(b) step 6), and "O-P" represents the remaining overhead from the interaction between the two phases (Figure 2(b) steps 1-2 and 4-5).
-"E2E" stands for end-to-end training latency.
 It can be observed that the use of dynamic heterogeneous strategies improves the performance of propagation.
-However, the interaction between optimization and propagation becomes slower due to the additional communication overhead introduced by LNK operators.
+However, the interaction between optimization and propagation becomes slower due to the additional communication overhead introduced by LNK.
 Despite this, the overhead remains relatively small and can be efficiently reduced by overlapping with computation, leading to only a minimal increase in latency.
 Moreover, this additional latency does not scale with the number of tokens per iteration, demonstrating the viability of disaggregating optimization and propagation.
-9 Related Work
-Optimizations of distributed training.
+9. Related Work
+Optimizations of distributed training
 Numerous studies aim to enhance the efficiency of distributed training.
-In the realm of parallel methods, advancements in pipeline parallel focus on minimizing bubble time through enhanced scheduling [69, 49, 39, 85], while tensor parallel has progressed by overlapping communication with computation [82, 37, 11, 40].
-In the context of model states sharding, efforts have been made to balance memory usage and latency through finer-grained sharding [13], efficient communication [81, 57], and multi-level storage solutions [73, 90].
+In the realm of parallel methods, advancements in pipeline parallel focus on minimizing bubble time through enhanced scheduling (Qi et al., 2023; Lamy-Poirier, 2023; Jiang et al., 2024a; Wu et al., 2024), while tensor parallel has progressed by overlapping communication with computation (Wang et al., 2023b; Jangda et al., 2022; Chen et al., 2024c; Jiang et al., 2024b).
+In the context of model states sharding, efforts have been made to balance memory usage and latency through finer-grained sharding (Chen et al., 2024a), efficient communication (Wang et al., 2023a; Maurya et al., 2024), and multi-level storage solutions (Ren et al., 2021; Zhao et al., 2025).
 Notably, these approaches are compatible with ours.
-There are also efforts focusing on automatically searching for the optimal parallel strategies [91, 38, 58, 84].
+There are also efforts focusing on automatically searching for the optimal parallel strategies (Zheng et al., 2022; Jia et al., 2019; Miao et al., 2022; Wang et al., 2024a; Li et al., 2025a; Wang et al., 2025b).
 However, they only consider homogeneous strategies and overlook the data imbalance issues.
-Resources disaggregation.
-The idea of disaggregating storage from compute resources has been well-studied in cloud computing [74, 31, 53] and utilized in many AI applications [3, 42].
-Similar concepts have also gained attention in large Transformer model inference, where disaggregating the prefilling and decoding phases allows for better throughput [67, 34, 92].
+Resources disaggregation
+The idea of disaggregating storage from compute resources has been well-studied in cloud computing (Shan et al., 2018; Guo et al., 2023; Lim et al., 2012) and utilized in many AI applications (Audibert et al., 2023; Jin et al., 2024).
+Similar concepts have also gained attention in large Transformer model inference, where disaggregating the prefilling and decoding phases allows for better throughput (Patel et al., 2024; Hu et al., 2024; Zhong et al., 2024).
 In our system, we adopt a similar design philosophy by separating the storage-heavy optimization phase from the compute-intensive propagation phase, enabling flexible execution and transitions of parallel strategies.
-Intermediate representations of parallel strategies.
-Graph-based intermediate representations are widely utilized in deep learning systems [1, 50, 41, 32].
-By analyzing and optimizing these graphs, models can discover enhanced partitioning strategies and parallel schemes.
-Hydraulis extends this concept by simultaneously deducing and expressing graphs for different strategies.
+Intermediate representations of parallel strategies
+Graph-based intermediate representations are widely utilized in deep learning systems (Abadi et al., 2016; Lattner et al., 2021; Jin et al., 2020; Hagedorn et al., 2023).
+By analyzing and optimizing the graphs, models can discover enhanced and specialized parallel strategies.
+Hydraulis extends this by simultaneously deducing and expressing graphs for different strategies.
 Using subgraphs, it distinctly highlights the unique communication patterns of each strategy.
 Besides, any graph-based optimization techniques can be applied to Hydraulis, allowing for further refinement for each strategy.
-10 Conclusion and Future Work
+10. Conclusion and Future Work
 In this work, we addressed the often-overlooked issues of data sampling and packing imbalances in existing training systems, and introduced Hydraulis, a novel system that leverages parallel computing and data management co-designs to optimize the training performance of large Transformer models.
-Empirical results show that Hydraulis can achieve up to a 2.66×2.$6$ increase in throughput compared to existing systems.
-As a possible future direction, we intend to expand our strategy space by integrating more parallel methods, as our system designs are fully compatible with more forms of parallelism.
-We also plan to tackle data imbalance challenges in diverse domains (e.g., in the video domain, the significant variation in sequence lengths has already attracted considerable attention [14, 55, 75]).
+We proposed training with dynamic heterogeneous parallel strategies to address the data sampling imbalance and a two-stage sequence assignment to tackle the data packing imbalance.
+Empirical results show that Hydraulis can achieve up to a $2.66\times$ increase in throughput compared to state-of-the-art training systems.
+As a possible future direction, we intend to expand our strategy space by integrating more parallel methods (e.g., expert parallel (Lepikhin et al., 2020; Liu et al., 2025)), as our system designs are fully compatible with more forms of parallelism.
+We also plan to tackle data imbalance challenges in diverse domains (e.g., in the video domain, the significant variation in sequence lengths has already attracted considerable attention (Chen et al., 2024b; Liu et al., 2024; Shang et al., 2024; Xia et al., 2025)).
 We believe our work can be adapted to a broader range of scenarios and applications.
+Acknowledgements.
+This work is supported by National Natural Science Foundation of China (U23B2048, 62402011), PKU-Tencent joint research Lab, and High-performance Computing Platform of Peking University.
+Fangcheng Fu and Bin Cui are the corresponding authors.
 Appendix A Data Imbalance
@@
 Subsequently, we conduct experiments to examine whether the quadratic time complexity of the attention mechanism impacts end-to-end performance.
-Figure 14: Curve of maximum sequence length across iterations. We sample 1.6M tokens from the CommonCrawl dataset per iteration and track the longest sequence.
-A.1 Larger Batch Size
+Figure 17. Curve of maximum sequence length across iterations. We sample 1.6M tokens from the CommonCrawl dataset per iteration and track the longest sequence.
+A.1. Larger Batch Size
 In our experiment, we use 64 GPUs, processing 100K tokens per iteration.
@@
 While this fluctuation is less pronounced compared to the scenario of using 100K tokens per iteration, it still indicates an apparent inter-iteration imbalance.
-A.2 End-to-end Influence of Attention
+A.2. End-to-end Influence of Attention
 To better demonstrate the impact of the quadratic time complexity of the attention mechanism on end-to-end training, we train the 7B model.
-Sequences are packed in mini-batches using the first-fit-decreasing bin packing algorithm [66], with a maximum length of 32K.
+Sequences are packed in mini-batches using the first-fit-decreasing bin packing algorithm (NVIDIA, 2023d), with a maximum length of 32K.
 Each mini-batch contains 100K tokens, and we manually ensure that they will form 4 packed sequences (if not, we just skip that iteration for this demonstration experiment).
@@
 This variation primarily stems from differences in the computational cost of the attention mechanism, while other computational operations and communication overhead remain relatively stable.
-Appendix B Cost Model
+Appendix B Communication in O-P Disaggregation
+B.1. Communication Bound
+In this section, we provide formal proof of the communication bound in O-P disaggregation.
+We consider $N$ GPUs and a model parameter of size $W$, and analyze the communication volume when each GPU participates as either a receiver or a sender during the pull operation, denoted as $V_{\text{recv}}$ and $V_{\text{send}}$, respectively.
+For the receiver side analysis, consider a GPU belonging to a pipeline that adopts a parallel scheme $P=\langle TP,PP,CP\rangle$ during the propagation phase.
+The GPU must receive the complete parameter partition corresponding to its pipeline stage.
+Specifically:
+  • If the parameter resides in the GPU’s current pipeline stage, it needs to receive a parameter partition of size $\frac{W}{TP}$.
+  • If the parameter does not belong to the current stage, the receiving volume becomes 0 as no pull operation is required.
+Therefore, we have:
+$V_{\text{recv}}=\frac{W}{TP}\quad\text{or}\quad V_{\text{recv}}=0$
+For the sender side analysis, our fully-sharded partitioning strategy during the optimization phase ensures each GPU stores exactly one parameter slice of size $\frac{W}{N}$.
+When $M$ GPUs (where $M\leq N$) request this particular parameter slice during propagation, the sending GPU’s communication volume is given by:
+$V_{\text{send}}=\frac{W}{N}\times M$
+Based on this, we can directly derive the worst-case communication scenario, which occurs when the parallel strategy takes the form $S=\langle 1,1,C\rangle\times\frac{N}{C}$, where $C$ is any positive integer divisor of $N$.
+This configuration represents the case where:
+  • The system contains no tensor or pipeline parallelism ($TP=1$, $PP=1$).
+  • Every GPU requires all parameter slices ($M=N$).
+Consequently, each GPU’s communication volume reaches its theoretical maximum:
+$V_{\text{recv}}=\frac{W}{1}=W\quad\text{and}\quad V_{\text{send}}=\frac{W}{N}\times N=W$
+This worst-case scenario exactly corresponds to training with a homogeneous fully-sharded ZeRO-1 strategy, where the communication cost per GPU is bounded by $W$ and becomes independent of the total GPU count $N$.
+An analogous analysis applies to the push operation.
+B.2. Computation-Communication Overlap
+In this section, we analyze the condition for overlapping communication with computation based on the derived communication bound.
+For a matrix multiplication operation under the pipeline with parallel scheme $\langle\mathit{TP},\mathit{PP},\mathit{CP}\rangle$, consider input $X$ with shape $t/\mathit{CP},h]$ and weight matrix $W$ with shape $h,h^{\prime}/\mathit{TP}]$, where $t$ denotes the number of tokens, $h$ and $h^{\prime}$ represent hidden state dimensions.
+During the computation of this matrix multiplication, we aim to pull the parameter for the next matrix multiplication.
+As established in Appendix B.1, the maximum amount of data communicated would be $h\times h^{\prime}$ (i.e., the whole parameter size).
+Assuming half-precision (e.g., BF16) is used, the total communication volume is:
+$\text{Comm}_{\text{bytes}}=2hh^{\prime}$
+And the computational workload (FLOPs) for the matrix multiply (counting multiply-add as 2 operations) is:
+$\text{FLOPs}=2\cdot\underbrace{\left(\frac{t}{\mathit{CP}}\right)}_{\text{rows}}\cdot\underbrace{h}_{\text{mid-dim}}\cdot\underbrace{\left(\frac{h^{\prime}}{\mathit{TP}}\right)}_{\text{cols}}=\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}}$
+Let $F$ denote the GPU’s half-precision compute capability (FLOPS) and $B$ the interconnect bandwidth (bytes/sec). The computation time $T_{\text{comp}}$ and communication time $T_{\text{comm}}$ are:
+$T_{\text{comp}}$ $=\frac{\text{FLOPs}}{F}=\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}\cdot F}$
+$T_{\text{comm}}$ $=\frac{\text{Comm}_{\text{bytes}}}{B}=\frac{2hh^{\prime}}{B}$
+To overlap communication with computation, we require:
+$T_{\text{comp}}\geq T_{\text{comm}}$
+Substituting the expressions and simplifying:
+$\frac{2thh^{\prime}}{\mathit{CP}\cdot\mathit{TP}\cdot F}\geq\frac{2hh^{\prime}}{B}$
+Canceling common terms ($2hh^{\prime}\neq 0$):
+$\frac{t}{\mathit{CP}\cdot\mathit{TP}\cdot F}\geq\frac{1}{B}$
+Solving for $t$:
+$t\geq\mathit{TP}\cdot\mathit{CP}\cdot\frac{F}{B}$
+This indicates that overlapping requires the token count $t$ to exceed a threshold determined solely by the parallel scheme ($\mathit{TP}$, $\mathit{CP}$) and hardware ratio ($F/B$), independent of parameter dimensions $h$, $h^{\prime}$.
+As a concrete example, consider an A800 cluster running the 32B model with a pipeline configured with $\left<16,1,1\right$>, corresponding to Figure 15(a).
+The hardware specifications are (assuming the worst-case scenario where all communication goes through InfiniBand):
+$F$ $=19.5\times 10^{12}\text{ FLOPS}$
+$B$ $=200\times 10^{9}\text{ bytes/sec (worst-case)}$
+The token threshold is then:
+$t_{\text{threshold}}$ $=\mathit{TP}\cdot\mathit{CP}\cdot\frac{F}{B}=16\times 1\times\frac{19.5\times 10^{12}}{200\times 10^{9}}=1560$
+This means that, theoretically, sequence lengths exceeding 1.5K after packing are sufficient to enable overlap (note that both GPU compute power and communication bandwidth may vary in real-world settings, so adjustments will be made based on empirical measurements).
+In practice, this condition is fully achievable in our setup.
+Appendix C Cost Model
 In this section, we elaborate on the memory model $\texttt{MaxLen}(P)$, which determines the maximum sequence length, and the latency model $\texttt{T}(l,P)$, which estimates the propagation time for a specific sequence length $l$, given a parallel scheme $P$.
-Figure 15: Latency breakdown and fluctuation when training LLaMA 7B on 8 Nvidia A800 GPUs.
+Figure 18.
+Latency breakdown and fluctuation when training LLaMA 7B on 8 Nvidia A800 GPUs.
 The left uses pipeline parallel (PP) without data parallel (DP), while the right uses data parallel (DP) without pipeline parallel (PP).
-B.1 Memory Model
+C.1. Memory Model
 Consider a model with hidden size $H$, number of layers $\overline{\textup{layers}}$, and vocab size $V$, and there are $N$ available GPUs.
-During the propagation phase, we utilize the parallel scheme $TP,PP\right>$, while the optimization phase employs a fully sharded strategy.
+During the propagation phase, we utilize the parallel scheme $\left<TP,PP,CP\right$>, while the optimization phase employs a fully sharded strategy.
 For a sequence length $l$ (i.e., micro-batch shape $[1,l]$), memory usage is calculated as follows.
-(a) For activations, the memory cost per micro-batch is $\frac{\overline{\textup{layers}}}{PP}\times\frac{l\times H}{TP}\times A$, where $A$ is a constant related to the model architecture, determined through profiling.
-The cumulative memory peak of all micro-batches (accumulated $PP$ times in the first pipeline) is $\overline{\textup{layers}}\times\frac{l\times H}{TP}\times A$.
-(b) For model states, given that some are fully sharded while others are related to $TP,PP\right>$, the total memory usage of all layers is $\frac{\overline{\textup{layers}}\times H^{2}}{N}\times B\times\alpha+\frac{\overline{\textup{layers}}}{PP}\times\frac{H^{2}}{TP}\times B\times\left(1-\alpha\right)$.
+(a) For activations, the memory cost per micro-batch is $\frac{\overline{\textup{layers}}}{PP}\times\frac{l\times H}{TP\times CP}\times A$, where $A$ is a constant related to the model architecture, determined through profiling.
+The cumulative memory peak of all micro-batches (accumulated $PP$ times in the first pipeline) is $\overline{\textup{layers}}\times\frac{l\times H}{TP\times CP}\times A$.
+(b) For model states, given that some are fully sharded while others are related to $\left<TP,PP,CP\right$>, the total memory usage of all layers is $\frac{\overline{\textup{layers}}\times H^{2}}{N}\times B\times\alpha+\frac{\overline{\textup{layers}}}{PP}\times\frac{H^{2}}{TP}\times B\times\left(1-\alpha\right)$.
 Here, $B$ is a constant related to the model structure, obtainable via profiling, and $\alpha$ is associated with the distribution and precision of model states.
@@
 The total memory peak is the sum of these costs, forming a linear function of $l$.
-Then, the x-coordinate of this function at the y-coordinate corresponding to GPU memory capacity minus a safety margin provides the maximum sequence length $\texttt{MaxLen}(P)$ supported by $P=\left<TP,PP\right>$.
+Then, the x-coordinate of this function at the y-coordinate corresponding to GPU memory capacity minus a safety margin provides the maximum sequence length $\texttt{MaxLen}(P)$ supported by $P=\left<TP,PP,CP\right$>.
 It’s important to note that we can apply a similar modeling approach when incorporating other forms of parallel methods, such as context parallel and sequence parallel.
-B.2 Latency Model
+C.2. Latency Model
 The latency is directly modeled as $\texttt{T}\left(l,P_{k}\right)=a_{k}\times l^{2}+b_{k}\times l+c_{k}$, where $a_{k},b_{k},c_{k}$ are coefficients related to $P_{k}$.
@@
 By profiling for $P_{k}$ and adjusting $l$, we fit the curve and determine these coefficients.
-Appendix C Enumeration Pruning
+Appendix D Enumeration Pruning
 (a) Efficiency comparison.
 (b) Bubble comparison.
-Figure 16: (a) LLaMA 13B training efficiency across different parallel schemes, measured in tokens per second.
-(b) Latency changes pre- and post-sequence packing for two equal-length sequences when $PP=2$, "F" and "B" indicate forward and backward propagation, respectively.
+Figure 19.
+(a) LLaMA 13B training efficiency across different parallel schemes, measured in tokens per second.
+(b) Latency changes pre- and post-sequence packing for two equal-length sequences when $PP=2$, “F” and “B” indicate forward and backward propagation, respectively.
 In this section, we will further discuss two trade-offs during packing as a supplement.
-Based on this, we will perform pruning optimization to avoid some unnecessary enumeration of the number of micro-batches.
-  1. 1.
+Based on this, we can perform pruning optimization to avoid some unnecessary enumeration of the number of micro-batches.
+  1. (1)
     Less packing, smaller micro-batch size, lower GPU utilization: As shown in Figure 19(a), we assess the GPU’s efficiency in processing sequences of specific lengths by dividing the propagation latency by the number of tokens in the sequence.
 Our observations indicate that efficiency approaches optimal only when the sequence length exceeds a certain threshold, which is related to intra-op parallel (e.g., TP) but independent of inter-op parallel (e.g., PP).
-This is due to GEMMs (General Matrix Multiplications) [64]: intra-op parallel reduces the computational load of single matrix multiplication, while inter-op parallel only reduces the number of matrix multiplications.
-When sequences are short or the intra-op parallel degree is large, the scale of a single matrix multiplication becomes small, leading to under-utilization of the GPU’s SMs and significant kernel launch overhead [63].
+This is due to GEMMs (General Matrix Multiplications) (NVIDIA, 2023b): intra-op parallel reduces the computational load of single matrix multiplication, while inter-op parallel only reduces the number of matrix multiplications.
+When sequences are short or the intra-op parallel degree is large, the scale of a single matrix multiplication becomes small, leading to under-utilization of the GPU’s SMs and significant kernel launch overhead (NVIDIA, 2023a).
 Therefore, insufficient packing on some parallel schemes may result in many under-utilized micro-batches, reducing the performance.
-  2. 2.
+  2. (2)
     More packing, smaller number of micro-batches, higher pipeline bubble rate: As shown in Figure 19(b), packing inevitably reduces the number of micro-batches, increasing the bubble rate.
-In the example shown, there is about a 0.25×0.25$\times$0.25 × performance slowdown.
+In the example shown, there is about a $0.25\times$ performance slowdown.
 Therefore, packing sequences to the memory limit or the mini-batch maximum is not necessarily optimal.
@@
 $V\in\left[\max\left(\frac{\sum\limits_{i=1}^{U}l_{i}}{\texttt{MaxLen}(P)},1\right),\min\left(\frac{\sum\limits_{i=1}^{U}l_{i}}{\texttt{UtilLen}(P)},U\right)\right]$
-Appendix D Details of Sequence Dispatching across Pipelines
+Appendix E Details of Sequence Dispatching across Pipelines
 In this section, we begin by deriving Eq.
@@
 (3)) and compare it to the precise solution.
-D.1 Derivation of the Lower Bound
+E.1. Derivation of the Lower Bound
 Input: Sequence lengths $\{l_{1},\ldots,l_{B}\}$. Maximum strategy indices $\{J_{1},\ldots,J_{B}\}$. Parallel schemes of the pipelines $\left\{{P}_{1},\ldots,{P}_{D}\right\}$. Number of trials $T$;
@@
 3
-4for trial $=1$ to $T$ do in parallel
+4for trial =1=1 to $T$ do in parallel
 5   Randomize permutation $\pi$ to shuffle sequences;
@@
 $\small\sum_{i=1}^{B}m_{i,d}\times\texttt{T}(l_{i},P_{d})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,d}\times l_{i}\right\},P_{d}\right)\times\left(\texttt{PP}\left(P_{d}\right)-1\right)$
-D.2 Approximate Solver
-[…]
-D.3 Performance Comparison
-Figure 17: Comparison of estimated propagation latency and actual end-to-end latency between greedy dispatching and ideal dispatching.
+E.2. Approximate Solver
+Our developed approximation algorithm is presented in Alg.
+1.
+We start by randomly shuffling the sequences and then dispatch them one by one (lines 1-6).
+For each dispatch, we aim to prevent the current slowest pipeline $P_{j}$ from worsening, by considering the accumulated sequence cost $C_{j}$ and extra cost $E_{j}$ (lines 7-14).
+Specifically, for each potential pipeline $j$ of the current sequence $i$ (line 7), we simulate dispatching the sequence to observe the cost change.
+We first compute the maximum sequence length based on the current sequence and those already dispatched to the pipeline (line 8).
+Then, we calculate the new accumulated cost $C_{j}^{\prime}$ by adding the latency of the current sequence $i$ on pipeline $j$ (line 9).
+After determining the new extra cost $E_{j}^{\prime}$, which is associated with the maximum sequence length and the pipeline parallel (PP) degree (line 10), we evaluate the maximum total cost $O_{\text{max}}$ across all pipelines (line 11).
+Afterward, we select the pipeline $j^{*}$ that minimizes $O_{\text{max}}$ and update the dispatch matrix and costs accordingly (lines 12-14).
+This process (i.e., a trial) is repeated $T$ times to select the best result (lines 15-17).
+[^8] TT is set to 100 in all our experiments.
+Since multiple trials can be parallelized using multiple processes, the actual execution speed is relatively fast (details in Appendix F), and the gap compared to the optimal solution of Eq.
+(3) is relatively small (details in Appendix E.3).
+E.3. Performance Comparison
+Figure 20.
+Comparison of estimated propagation latency and actual end-to-end latency between greedy dispatching and ideal dispatching.
 The experiments use a 32B model on the CommonCrawl dataset, with 64 GPUs.
 Each iteration handles 200,000 tokens with a maximum sequence length of 32K.
+Table 2.
+Solving time for each subproblem.
+The time cost for ideal dispatching, greedy dispatching, and packing is listed for 10 iterations combined.
+Hydraulis ultimately chooses the greedy dispatching scheme over the ideal dispatching scheme due to its faster speed.
 Strategy Proposal Ideal Dispatching Greedy Dispatching Packing
 64 GPUs & 200K tokens per iteration 2.03s 40.21s 5.46s 10.15s
-1024 GPUs & 3.2M tokens per iteration 107.21s >1000s 38.56s 13.46s
-Table 2: Solving time for each subproblem.
-The time cost for ideal dispatching, greedy dispatching, and packing is listed for 10 iterations combined.
-Hydraulis ultimately chooses the greedy dispatching scheme over the ideal dispatching scheme due to its faster speed.
+1024 GPUs & 3.2M tokens per iteration 107.21s ¿1000s 38.56s 13.46s
+Table 3.
+Optimal training configurations.
+DP refers to the data parallel degree, TP to the tensor parallel degree, $P$ to the pipeline parallel degree, CP to the context parallel degree, SP to the Ulysses sequence parallel degree, and AC indicates the use of activation checkpointing.
+For HotSPa, there are 2-3 distinct strategies.
 LLaMA 7B 8 GPUs 16 GPUs
@@
 DeepSpeed DP8SP4 + AC DP4SP8 + AC DP16SP4 + AC DP8SP8 + AC
-Megatron-LM DP2TP8PP2 DP2TP16 DP4TP8PP2 DP4TP16
+Megatron-LM DP2TP8PP2 DP2TP8CP2 DP4TP8PP2 DP4TP8CP2
 HotSPa DP2TP8PP2, DP4TP4PP2 DP2TP16, DP2TP8PP2, DP4TP4PP2 DP4TP8PP2, DP8TP4PP2 DP4TP16, DP4TP8PP2, DP8TP4PP2
-Table 3: Optimal training configurations.
-DP refers to the data parallel degree, TP to the tensor parallel degree, PP to the pipeline parallel degree, SP to the Ulysses sequence parallel degree, and AC indicates the use of activation checkpointing.
-For HotSPa, there are 2-3 distinct strategies.
+Table 4. HotSPa’s switching overhead and iteration time comparison.
+Configuration Tokens per Iteration HotSPa Switching Time HotSPa Iteration Time (incl. Switching) Hydraulis Iteration Time
+(a) 7B, 32 GPUs 100 1.54s 5.86s 2.21s
+(b) 32B, 32 GPUs 100 2.17s 12.21s 7.21s
+(c) 32B, 64 GPUs 200 2.21s 13.19s 7.22s
 We analyze the gap between our proposed dispatching algorithm, Alg.
@@
 This indicates that there is no significant performance difference between the two approaches.
-Considering the subsequent time analysis in Appendix E, we ultimately opt for greedy dispatching due to its faster solution speed and superior scalability.
-Appendix E Time Cost and Scalability
+Considering the subsequent time analysis in Appendix F, we ultimately opt for greedy dispatching due to its faster solution speed and superior scalability.
+Appendix F Time Cost and Scalability
 In this section, we discuss the computation times of all mathematical models and their scalability.
@@
 This design ensures that GPUs remain fully saturated with training workloads throughout execution, thereby achieving efficient utilization of computational resources.
-Appendix F Optimal Training Configurations
-Table 3 illustrates the manually tuned configurations that achieve the best training performance under different scenarios in § 8.2.
+Appendix G Optimal Training Configurations
+Table 3 illustrates the manually tuned configurations that achieve the best training performance under different scenarios in §8.2. In terms of HotSPa:
   • For scenario F1, with a maximum sequence length of 16K, two strategies are employed: (a) The first strategy handles sequences of 4K to 16K. (b) The second strategy handles sequences less than 4K.
@@
 (c) The third strategy handles sequences less than 4K.
+Appendix H HotSPa’s Switching Overhead
+In this section, we analyze the switching overhead of HotSPa under three configurations, as detailed in Table 4.
+From configuration (a) to (b), we scale the model size, and from (b) to (c), we scale the total training setup, proportionally to LLaMA’s configuration (16M tokens per iteration for 8192 GPUs).
+First, when the training workload is relatively small, HotSPa’s switching overhead becomes significant.
+For example, in configuration (a), 1.54s out of the total 5.86s per iteration (26%) is spent on strategy switching, negatively impacting end-to-end performance.
+Second, when scaling from configuration (b) to (c), doubling the data parallel (DP) degrees causes HotSPa’s switching to occur independently within each 32-GPU group.
+This keeps the switching overhead relatively stable.
+However, since the tokens per iteration also scale, the overhead remains consistently around 17% of the total iteration time.
+(Note that our experimental environment is equipped with 400 GB/s NVLink and 200 GB/s IB, which are sufficiently fast.) In contrast, Hydraulis overlaps strategy switching with computation via O-P disaggregation (as evaluated in §8, Table 1), making it a more scalable solution.
+Finally, even if the switching overhead were excluded, HotSPa would still lag behind Hydraulis.
+This is due to the advantages of Hydraulis in terms of strategy optimization and sequence assignment: For one thing, HotSPa only supports homogeneous strategies at any given time, whereas Hydraulis enables dynamic heterogeneous strategies.
+For another, Hydraulis incorporates fine-grained sequence dispatching and packing, which HotSPa lacks.